### v1

# Nonparametric regression using over-parameterized shallow ReLU neural networks

Abstract
It is shown that over-parameterized neural networks can achieve minimax optimal rates of convergence (up to logarithmic factors) for learning functions from certain smooth function classes, if the weights are suitably constrained or regularized. Specifically, we consider the nonparametric regression of estimating an unknown $d$-variate function by using shallow ReLU neural networks. It is assumed that the regression function is from the Hölder space with smoothness $\alpha<(d+3)/2$ or a variation space corresponding to shallow neural networks, which can be viewed as an infinitely wide neural network. In this setting, we prove that least squares estimators based on shallow neural networks with certain norm constraints on the weights are minimax optimal, if the network width is sufficiently large. As a byproduct, we derive a new size-independent bound for the local Rademacher complexity of shallow ReLU neural networks, which may be of independent interest.
Keywords: neural networks, nonparametric regression, over-parameterization, rate of convergence
Mathematics Subject Classification: 62G08, 68T07
1 Introduction
In nonparametric regression, we are given a data set of $n$ samples $\mathcal{D}_{n}=\{(X_{i},Y_{i})\}_{i=1}^{n}$, which are independent and identically distributed as a $\mathbb{R}^{d}\times\mathbb{R}$-valued random vector $(X,Y)$. The statistical problem is to estimate the so-called regression function $h(x)=\mathbb{E}[Y|X=x]$ from the observed data $\mathcal{D}_{n}$. One of the most popular estimators is the least squares
$\widehat{f}_{n}\in\operatorname*{argmin}_{f\in\mathcal{F}}\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})-Y_{i})^{2},$
where $\mathcal{F}$ is a suitably chosen function class. In recent years, due to the breakthrough of deep learning, there is an increasing interest in the literature in studying the performance of least squares when $\mathcal{F}$ is parameterized by neural networks [Bauer and Kohler, 2019; Chen et al., 2022; Jiao et al., 2021; Kohler and Langer, 2021; Nakada and Imaizumi, 2020; Schmidt-Hieber, 2020; Suzuki, 2019; Yang and Zhou, 2023].
To analyze the rate of convergence for the least squares estimator $\widehat{f}_{n}$, one often assume that the regression function $h$ is in certain smooth function class $\mathcal{H}$. Typically, we can bound the error of $\widehat{f}_{n}$ as
$\mathbb{E}_{X}\left[\left|\widehat{f}_{n}(X)-h(X)\right|^{2}\right]\lesssim\mathcal{E}_{app}(\mathcal{H},\mathcal{F})+\mathcal{C}_{n}(\mathcal{F}),$ (1.1)
where, for two quantities $A$ and $B$, $A\lesssim B$ (or $B\gtrsim A$) denotes the statement that $A\leq cB$ for some constant $c>0$ (we will also denote $A\asymp B$ when $A\lesssim B\lesssim A$). In the error decomposition (1.1), the approximation error $\mathcal{E}_{app}(\mathcal{H},\mathcal{F})$ quantifies how well we can approximate $h\in\mathcal{H}$ by functions in $\mathcal{F}$, and the sample complexity $\mathcal{C}_{n}(\mathcal{F})$ measures how well the learned function $\widehat{f}_{n}$ generalizes to unseen data. For example, the complexity $\mathcal{C}_{n}(\mathcal{F})$ is often controlled by the covering number, VC dimension [Vapnik and Chervonenkis, 1971] or Rademacher complexity [Bartlett and Mendelson, 2002] of $\mathcal{F}$. In most analysis for neural networks, the approximation error and the complexity are bounded by the size of the networks, e.g. the number of parameters or neurons. In particular, when $\mathcal{H}=\mathcal{H}^{\alpha}$ is the unit ball of Hölder space with smoothness index $\alpha>0$, the results of [Schmidt-Hieber, 2020; Kohler and Langer, 2021] established the rate $n^{-\frac{2\alpha}{d+2\alpha}}$ for deep neural networks, which is minimax optimal [Stone, 1982]. However, in these analyses, one has to use under-parameterized neural networks (the number of parameters is less than the sample size $n$) in order to get the optimal trade-offs between $\mathcal{E}_{app}(\mathcal{H},\mathcal{F})$ and $\mathcal{C}_{n}(\mathcal{F})$. But, in practical applications of deep neural networks, the number of parameters is often much larger than the number of samples. Hence, these analyses can not fully explain the success of neural networks used in practice. To obtain rates of convergence in the over-parameterized regime, the recent works [Jiao et al., 2023; Yang and Zhou, 2023] proposed to bound $\mathcal{E}_{app}(\mathcal{H},\mathcal{F})$ and $\mathcal{C}_{n}(\mathcal{F})$ by the size of the weights, i.e. certain norms of the weight matrices, rather than the size of networks. Using this idea, in [Yang and Zhou, 2023], we obtained nearly optimal rates for over-parameterized neural networks.
The purpose of this paper is to show that over-parameterized neural networks can achieve the minimax optimal rates for estimating certain smooth functions, if the weights are suitably constrained or regularized. Specifically, we consider shallow neural networks with ReLU activation
$f_{\theta}(x):=\sum_{i=1}^{N}a_{i}\sigma((x^{\intercal},1)w_{i}),$
where $\theta:=(a_{1},w_{1}^{\intercal},\dots,a_{N},w_{N}^{\intercal})^{\intercal}$ denotes the vector of trainable parameters and $\sigma(t):=\max\{t,0\}$ is the rectified linear unit (ReLU). Following the ideas in [Jiao et al., 2023; Yang and Zhou, 2023], we constrain the weights by $\kappa(\theta)=\sum_{i=1}^{N}|a_{i}|\|w_{i}\|_{2}\leq M$ and study the performance of the constrained least squares
$\operatorname*{argmin}_{\kappa(\theta)\leq M}\frac{1}{n}\sum_{i=1}^{n}(f_{\theta}(X_{i})-Y_{i})^{2}.$
As in [Yang and Zhou, 2023], we assume that the regression function $h$ belongs to the Hölder class $\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$ or the unit ball of the variation space $\mathcal{F}_{\sigma}(1)$, which can be regarded as an infinitely wide neural network with weight constraint $\kappa(\theta)\leq 1$ (see Section 2 for details). The paper [Yang and Zhou, 2023] proved the rates $n^{-\frac{2\alpha}{d+3+2\alpha}}$ for $\mathcal{H}^{\alpha}$ and $n^{-1/2}$ for $\mathcal{F}_{\sigma}(1)$, if $M$ is chosen properly and the network width $N$ is sufficiently large (i.e. constrained over-parameterized networks). We improve these rates to $n^{-\frac{2\alpha}{d+2\alpha}}$ and $n^{-\frac{d+3}{2d+3}}$, which are minimax optimal for $\mathcal{H}^{\alpha}$ and $\mathcal{F}_{\sigma}(1)$ respectively. The improvement is from the use of localization technique [Bartlett et al., 2005; Koltchinskii, 2006] in estimating the complexity $\mathcal{C}_{n}(\mathcal{F})$ of neural networks. One of our main technical contributions is a new bound for the local Rademacher complexity of shallow neural networks (see Definition 3.2 and Theorem 3.3):
$\mathcal{R}_{n}(\mathcal{NN}(N,M);\delta)\lesssim\frac{\delta^{\frac{3}{d+3}}M^{\frac{d}{d+3}}}{\sqrt{n}}\sqrt{\log(nM/\delta)},$
where $\mathcal{NN}(N,M)$ is the set of neural network functions $f_{\theta}$ with width at most $N$ and constraint $\kappa(\theta)\leq M$. Our bound generalizes the the bound $\mathcal{O}(M/\sqrt{n})$ for Rademacher complexity in [Golowich et al., 2020], which is used in [Yang and Zhou, 2023] to control the sample complexity.
Besides the constrained least squares, we also study the corresponding regularized least squares
$\operatorname*{argmin}_{\theta}\frac{1}{n}\sum_{i=1}^{n}(f_{\theta}(X_{i})-Y_{i})^{2}+\lambda\kappa(\theta),$
where $\lambda>0$ is a tunable parameter. It is proven that this optimization problem has essentially the same solutions as the problem
$\operatorname*{argmin}_{\theta}\frac{1}{n}\sum_{i=1}^{n}(f_{\theta}(X_{i})-Y_{i})^{2}+\frac{\lambda}{2}\|\theta\|_{2}^{2}.$
Our second main result shows that the solutions of these two optimization problems also achieve the minimax optimal rates of learning functions in $\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$ and $\mathcal{F}_{\sigma}(1)$, if $\lambda$ is chosen properly. This gives a theoretical guarantee for the use of some regularization methods, such as [Neyshabur et al., 2015a].
The rest of the paper is organized as follows. In Section 2, we introduce some function classes of shallow neural networks and their approximation results. Section 3 presents our results on the rates of convergence for over-parameterized neural networks. In Section 4, we give some discussions on our results and related works. The proofs of our main theorems are given in Section 5.
2 The variation space of shallow neural networks
Let $\mathbb{B}^{d}=\{x\in\mathbb{R}^{d}:\|x\|_{2}\leq 1\}$ be the unit ball of $\mathbb{R}^{d}$, and $\mathbb{S}^{d}=\{x\in\mathbb{R}^{d+1}:\|x\|_{2}=1\}$ be the unit sphere of $\mathbb{R}^{d+1}$. We denote the space of signed Radon measures on $\mathbb{S}^{d}$ by $\mathcal{M}(\mathbb{S}^{d})$, and denote the total variation norm of $\mu\in\mathcal{M}(\mathbb{S}^{d})$ by $\|\mu\|=|\mu|(\mathbb{S}^{d})$. It is well known that $\mathcal{M}(\mathbb{S}^{d})$ is the dual space of $C(\mathbb{S}^{d})$ [Folland, 1999, Section 7.3]. We are interested in the function class $\mathcal{F}_{\sigma}$ that contains continuous functions $f_{\mu}:\mathbb{B}^{d}\to\mathbb{R}$ of the integral form
[…]
where $\sigma$ is the ReLU function. Note that the integral representation is not unique for $f\in\mathcal{F}_{\sigma}$. For example, if $\mu$ is supported on the set
$S_{-}:=\left\{(x_{1},\dots,x_{d+1})^{\intercal}\in\mathbb{S}^{d}:x_{d+1}\leq-\sqrt{2}/2\right\},$ (2.2)
then $f_{\mu}=0$ is the zero function on $\mathbb{B}^{d}$. Following Bach [2017] and Yang and Zhou [2023], we can define a norm on $\mathcal{F}_{\sigma}$ by
$\gamma(f):=\inf\left\{\|\mu\|:f=f_{\mu},\mu\in\mathcal{M}(\mathbb{S}^{d})\right\}.$
By the compactness of closed balls in $\mathcal{M}(\mathbb{S}^{d})$ (due to Prokhorov’s theorem), the infimum defining $\gamma(f)$ is attained by some measure $\mu$. One can show that $\mathcal{F}_{\sigma}$ equipped with the norm $\gamma$ is a Banach space. Since the norm is defined through the variation norm, we call $\mathcal{F}_{\sigma}$ the variation space (of shallow ReLU neural networks). For any $M>0$, we denote
$\mathcal{F}_{\sigma}(M):=\{f\in\mathcal{F}_{\sigma}:\gamma(f)\leq M\}.$
This function class can be thought of as an infinitely wide neural network with a constraint on its weights. There are several other definitions and characterizations of function spaces corresponding to infinitely wide neural networks in recent studies [Bartolucci et al., 2023; Ongie et al., 2020; Parhi and Nowak, 2022, 2023; Savarese et al., 2019; Siegel and Xu, 2020, 2022, 2023]. We will give more discussion in Section 4.
The neural networks with finite neurons are corresponding to discrete measures $\mu\in\mathcal{M}(\mathbb{S}^{d})$ that support on finitely many points. Let us denote the collection of these measures by $\mathcal{M}_{disc}(\mathbb{S}^{d})$, which is a linear subspace of $\mathcal{M}(\mathbb{S}^{d})$, and define the corresponding function class $\mathcal{F}_{\sigma,disc}:=\{f_{\mu},\mu\in\mathcal{M}_{disc}(\mathbb{S}^{d})\}$. Similar to the norm $\gamma$ defined on $\mathcal{F}_{\sigma}$, we can also define a norm on $\mathcal{F}_{\sigma,disc}$ by
$\kappa(f):=\inf\left\{\|\mu\|:f=f_{\mu},\mu\in\mathcal{M}_{disc}(\mathbb{S}^{d})\right\}.$
By definition, $\gamma(f)\leq\kappa(f)$ for all $f\in\mathcal{F}_{\sigma,disc}$. But we do not know whether these two norms are the same for $f\in\mathcal{F}_{\sigma,disc}$. Nevertheless, it is good enough to bound $\gamma(f)$ by $\kappa(f)$ in our theory. We will see soon that the norm $\kappa(f)$ can be estimated in practice, and hence it can be used in statistical estimations.
For any $N\in\mathbb{N}$, the function class $\mathcal{NN}(N)\subseteq\mathcal{F}_{\sigma,disc}$ of shallow neural networks with (at most) $N$ neurons can be parameterized in the form
$f_{\theta}(x):=\sum_{i=1}^{N}a_{i}\sigma((x^{\intercal},1)w_{i}),\quad a_{i}\in\mathbb{R},w_{i}\in\mathbb{R}^{d+1},x\in\mathbb{B}^{d},$ (2.3)
where $\theta$ denote the vector of parameters
$\theta:=(a_{1},w_{1}^{\intercal},\dots,a_{N},w_{N}^{\intercal})^{\intercal}\in\mathbb{R}^{(d+2)N}.$
Note that, in this parameterization, we do not restrict the internal weights $w_{i}$ to $\mathbb{S}^{d}$. Thanks to the homogeneity of ReLU function, it is easy to see that $f_{\theta}=f_{\mu}$ with discrete measure $\mu=\sum_{i=1}^{N}c_{i}\delta_{v_{i}}$, where $v_{i}=w_{i}/\|w_{i}\|_{2}$ and $c_{i}=a_{i}\|w_{i}\|_{2}$ if $\|w_{i}\|_{2}\neq 0$, and $c_{i}=0$ if $\|w_{i}\|_{2}=0$. This implies that we can estimate $\kappa(f_{\theta})$ by the following quantity
$\kappa(\theta):=\sum_{i=1}^{N}|a_{i}|\|w_{i}\|_{2}.$
With a slight abuse of notation, we have denoted it by the same notation as the norm $\kappa$. The reason is given by the following theorem, which shows that $\kappa(f_{\theta})$ is equivalent to the minimum of $\kappa(\theta)$ over all parameterization (2.3).
Theorem 2.1.
For any $f\in\mathcal{NN}(N)$,
$\kappa(f)\leq\inf\left\{\kappa(\theta):f=f_{\theta},\theta\in\mathbb{R}^{(d+2)N}\right\}\leq 3\kappa(f).$
The proof is given in Section 5.1. In the proof, we actually show that any $f\in\mathcal{NN}(N)$ can be reduced to the following form
$f(x)=\sum_{k=1}^{K}c_{k}\sigma((x^{\intercal},1)v_{k})+(x^{\intercal},1)w,$ (2.4)
where $0\leq K\leq N$, $w\in\mathbb{R}^{d+1}$ and $v_{1},\dots,v_{K}\in S_{0}$ satisfy $v_{i}\neq\pm v_{k}$ for $1\leq i\neq k\leq K$. The set $S_{0}$ is defined by $S_{0}:=\mathbb{S}^{d}\setminus(S_{-}\cup S_{+})$, where $S_{-}$ is defined by (2.2) and $S_{+}:=\{-v:v\in S_{-}\}$. Note that the set $S_{0}$ contains vectors $v\in\mathbb{S}^{d}$ such that the function $x\mapsto\sigma((x^{\intercal},1)v)$ is nonlinear on $\mathbb{B}^{d}$. We prove that, for the reduced parameterization (2.4),
$\|w\|_{2}\lor\sum_{k=1}^{K}|c_{k}|\leq\kappa(f)\leq 2\|w\|_{2}+\sum_{k=1}^{K}|c_{k}|,$ (2.5)
where we use the notation $a\lor b:=\max\{a,b\}$. This gives a way to estimate $\kappa(f)$ in practice. Roughly speaking, the norm $\kappa(f)$ is a sum of the variation of the nonlinear part of $f$ and the norm of the linear part.
Now, we are ready to introduce the central object of this paper
$\mathcal{NN}(N,M):=\left\{f_{\theta}\in\mathcal{NN}(N):\kappa(\theta)\leq M\right\}.$
Since $\gamma(f_{\theta})\leq\kappa(f_{\theta})\leq\kappa(\theta)$, we have $\mathcal{NN}(N,M)\subseteq\mathcal{F}_{\sigma}(M)$. It was shown in [Yang and Zhou, 2023, Proposition 2.2] that $\mathcal{F}_{\sigma}(M)$ is the closure of $\cup_{N\in\mathbb{N}}\mathcal{NN}(N,M)$ in $C(\mathbb{B}^{d})$. The rates of approximation by $\mathcal{NN}(N,M)$ for smooth function classes are also studied in many works [Bach, 2017; Klusowski and Barron, 2018; Siegel and Xu, 2022; Yang and Zhou, 2023]. To describe these approximation results, let us first recall the classical notion of smoothness of functions. Given a smoothness index $\alpha>0$, we write $\alpha=r+\beta$ where $r\in\mathbb{N}_{0}:=\mathbb{N}\cup\{0\}$ and $\beta\in(0,1]$. Let $C^{r,\beta}(\mathbb{R}^{d})$ be the Hölder space with the norm
$\|f\|_{C^{r,\beta}(\mathbb{R}^{d})}:=\max\left\{\|f\|_{C^{r}(\mathbb{R}^{d})},\max_{\|s\|_{1}=r}|\partial^{s}f|_{C^{0,\beta}(\mathbb{R}^{d})}\right\},$
where $s=(s_{1},\dots,s_{d})\in\mathbb{N}_{0}^{d}$ is an multi-index and
$\|f\|_{C^{r}(\mathbb{R}^{d})}:=\max_{\|s\|_{1}\leq r}\|\partial^{s}f\|_{L^{\infty}(\mathbb{R}^{d})},\quad|f|_{C^{0,\beta}(\mathbb{R}^{d})}:=\sup_{x\neq y\in\mathbb{R}^{d}}\frac{|f(x)-f(y)|}{\|x-y\|_{2}^{\beta}},$
where we use $\|\cdot\|_{L^{\infty}}$ to denote the supremum norm, since we only consider continuous functions in this paper. We write $C^{r,\beta}(\mathbb{B}^{d})$ for the Banach space of all restrictions to $\mathbb{B}^{d}$ of functions in $C^{r,\beta}(\mathbb{R}^{d})$. The norm is given by $\|f\|_{C^{r,\beta}(\mathbb{B}^{d})}=\inf\{\|g\|_{C^{r,\beta}(\mathbb{R}^{d})}:g\in C^{r,\beta}(\mathbb{R}^{d})\mbox{ and }g=f\mbox{ on }\mathbb{B}^{d}\}$. For convenience, we will denote the unit ball of $C^{r,\beta}(\mathbb{B}^{d})$ by
$\mathcal{H}^{\alpha}:=\left\{f\in C^{r,\beta}(\mathbb{B}^{d}):\|f\|_{C^{r,\beta}(\mathbb{B}^{d})}\leq 1\right\}.$
Note that, for $\alpha=1$, $\mathcal{H}^{\alpha}$ is Lipschitz continuous function class.
In [Yang and Zhou, 2023], we characterized how well one can use neural networks in $\mathcal{F}_{\sigma}(M)$ to approximate functions in $\mathcal{H}^{\alpha}$. It was shown that if $\alpha>(d+3)/2$ then $\mathcal{H}^{\alpha}\subseteq\mathcal{F}_{\sigma}(M)$ for some constant $M$, and if $\alpha<(d+3)/2$ then
$\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{F}_{\sigma}(M)}\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim M^{-\frac{2\alpha}{d+3-2\alpha}}.$
For the critical value $\alpha=(d+3)/2$, it seems that the inclusion $\mathcal{H}^{\alpha}\subseteq\mathcal{F}_{\sigma}(M)$ also holds for some constant $M$, but [Yang and Zhou, 2023] only obtained the exponential approximation rate $\mathcal{O}(\exp(-\alpha M^{2}))$. Combining these results with the rate of approximation for $\mathcal{F}_{\sigma}(M)$ by its subclass $\mathcal{F}_{\sigma}(N,M)$ derived in [Bach, 2017; Siegel and Xu, 2022], one can obtain the following approximation bounds for shallow neural networks. Recall that we use the notation $a\lor b:=\max\{a,b\}$.
Theorem 2.2 ([Yang and Zhou, 2023]).
Let $d\geq 4$.
  1. (1)
    For $\mathcal{H}^{\alpha}$ with $20<\alpha<(d+3)/$ 2, we have
    $\sup_{h\in\mathcal{H}^{\alpha}}\inf_{f\in\mathcal{NN}(N,M)}\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim N^{-\frac{\alpha}{d}}\lor M^{-\frac{2\alpha}{d+3-2\alpha}}.$
  2. (2)
    For $\mathcal{F}_{\sigma}(1)$, there exists a constant $M\geq 1$ such that
    $\sup_{h\in\mathcal{F}_{\sigma}(1)}\inf_{f\in\mathcal{NN}(N,M)}\|h-f\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim N^{-\frac{d+3}{2d}}.$
As discussed by Yang and Zhou [2023], for $\mathcal{H}^{\alpha}$ with $20<\alpha<(d+3)/$ 2, the approximation rate is optimal in terms of the number of neurons $N$ and the norm constraint $M$. We will use this result to establish the optimal rate of convergence for nonparametric regression using shallow neural networks.
3 Main results
We consider the problem of nonparametric regression in classical setting. Suppose we have a data set of $n\geq 2$ samples $\mathcal{D}_{n}=\{(X_{i},Y_{i})\}_{i=1}^{n}$, which are independently and identically generated from a probability distribution supported on $\mathbb{B}^{d}\times\mathbb{R}$. Let $\mu$ be the marginal distribution of the covariate $X$ and $h(x):=\mathbb{E}[Y|X=x]$ denote the regression function. In this paper, we assume that $|Y|\leq B$ is bounded by some fixed constant $B>0$ and $h\in\mathcal{H}$, where we will study the cases that $\mathcal{H}$ is the unit ball of the variation space $\mathcal{F}_{\sigma}(1)$ or the Hölder class $\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$. Note that $\mathcal{H}^{\alpha}$ with $\alpha>(d+3)/2$ is included in the first case by scaling, because $\mathcal{H}^{\alpha}\subseteq\mathcal{F}_{\sigma}(M)$ for some $M$ [Yang and Zhou, 2023]. Let us denote the noises by
$\eta_{i}:=Y_{i}-h(X_{i}),\quad i=1,\dots,n,$
then $\mathbb{E}[\eta_{i}]=0$ and $|\eta_{i}|\leq 2B$. We will often denote the sample data points by the sequence $X_{1:n}:=(X_{1},\dots,X_{n})$ (similarly $\eta_{1:n}:=(\eta_{1},\dots,\eta_{n})$ denotes the sequence of noises). The empirical distribution is denoted by $\mu_{n}:=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}}$, and the associated $L^{2}(\mu_{n})$ norm is given by $\|f\|_{L^{2}(\mu_{n})}^{2}:=\frac{1}{n}\sum_{i=1}^{n}f(X_{i})^{2}$.
3.1 Least squares
One popular way to estimate the regression function $h\in\mathcal{H}$ is by constrained least squares
$\widehat{f}_{n}\in\operatorname*{argmin}_{f\in\mathcal{F}_{n}}\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})-Y_{i})^{2},$ (3.1)
where we consider the case that $\mathcal{F}_{n}=\mathcal{NN}(N_{n},M_{n})$ is parameterized by a shallow ReLU neural network. For simplicity, we assume here and in the sequel that the minimum above indeed exists. The performance of the estimation is measured by the expected risk
$\mathcal{L}(f):=\mathbb{E}_{(X,Y)}[(f(X)-Y)^{2}].$
It is equivalent to evaluate the estimator by the excess risk
$\|f-h\|_{L^{2}(\mu)}^{2}=\mathcal{L}(f)-\mathcal{L}(h).$
In the statistical analysis of learning algorithms, we often require that the hypothesis class is uniformly bounded. We define the truncation operator $\mathcal{T}_{b}$ with level $b>0$ for real-valued functions $f$ as
$\mathcal{T}_{b}f(x):=\begin{cases}f(x),&\quad\mbox{if }|f(x)|\leq b,\\ \,{\rm sgn}\,(f(x))b,&\quad\mbox{if }|f(x)|>b.\end{cases}$
It is easy to see that, if $b\geq\sup_{h\in\mathcal{H}}\|h\|_{L^{\infty}(\mathbb{B}^{d})}$, then $\|\mathcal{T}_{b}\widehat{f}_{n}-h\|_{L^{2}(\mu)}\leq\|\widehat{f}_{n}-h\|_{L^{2}(\mu)}$ for any $h\in\mathcal{H}$. Hence, truncating the output of the estimator $\widehat{f}_{n}$ appropriately dose not increase the excess risk. In the following, we will simply take $b=B$.
Theorem 3.1.
Let $d\geq 4$ and $\widehat{f}_{n}$ be the estimator (3.1) with $\mathcal{F}_{n}=\mathcal{NN}(N_{n},M_{n})$.
  1. (1)
    If $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$, we choose
    $M_{n}\asymp n^{\frac{d+3-2\alpha}{2d+4\alpha}},\quad N_{n}\gtrsim n^{\frac{d}{d+2\alpha}},$
    then the bound
    $\|\mathcal{T}_{B}\widehat{f}_{n}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{2\alpha}{d+2\alpha}}\log n$
    holds with probability at least $1-c_{1}\exp(-c_{2}n^{\frac{d}{d+2\alpha}}\log n)$ for some constants $c_{1},c_{2}>0$.
  2. (2)
    If $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, then there exist a constant $M\geq 1$ such that for any
    $M_{n}=M,\quad N_{n}\gtrsim n^{\frac{d}{2d+3}},$
    the bound
    $\|\mathcal{T}_{B}\widehat{f}_{n}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{d+3}{2d+3}}\log n$
    holds with probability at least $1-c_{1}\exp(-c_{2}n^{\frac{d}{2d+3}}\log n)$ for some constants $c_{1},c_{2}>0$.
Note that the (implied) constants in the theorem depend on $d$, $\alpha$ and $B$. Notice that we can always take $N_{n}\gtrsim n$, which gives justification for the use of over-parameterized neural networks in practice. The result also holds for $\mathcal{F}_{n}=\mathcal{F}_{\sigma}(M_{n})$, which is corresponding to $N_{n}=\infty$. The required minimal widths $N_{n}$ of neural networks match the result of [Yang and Zhou, 2023, Theorem 4.2] in under-parameterized regime. For over-parameterized shallow neural networks, Theorem 3.1 improves the rates of convergence in [Yang and Zhou, 2023, Theorem 4.7]. Ignoring logarithmic factors, the rates $n^{-\frac{2\alpha}{d+2\alpha}}$ for $\mathcal{H}^{\alpha}$ and $n^{-\frac{d+3}{2d+3}}$ for $\mathcal{F}_{\sigma}(1)$ are minimax optimal, as shown by [Stone, 1982; Parhi and Nowak, 2023; Yang and Zhou, 2023].
The proof of Theorem 3.1 is deferred to Section 5.2. Similar to the classical analysis of nonparametric regression [Györfi et al., 2002; Wainwright, 2019], our proof decomposes the error $\|\mathcal{T}_{B}\widehat{f}_{n}-h\|_{L^{2}(\mu)}^{2}$ as (1.1) into two terms: the approximation error $\mathcal{E}_{app}(\mathcal{H},\mathcal{F}_{n})$ and the complexity $\mathcal{C}_{n}(\mathcal{F}_{n})$. The approximation error can be handled by Theorem 2.2. We control the complexity by using the localization technique [Bartlett et al., 2005; Koltchinskii, 2006].
Definition 3.2 (Local complexity).
Let $\xi_{1:n}$ be a sequence of independent zero-mean random variables. For a given radius $\delta>0$ and a sequence of sample points $X_{1:n}$, we define the local complexity of a function class $\mathcal{F}$ at scale $\delta$ with respect to $\xi_{1:n}$ by
$\mathcal{G}_{n}(\mathcal{F};\delta,\xi_{1:n}):=\mathbb{E}_{\xi_{1:n}}\left[\sup_{f\in\mathcal{F},\|f\|_{L^{2}(\mu_{n})}\leq\delta}\left|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}f(X_{i})\right|\right].$
If each $\xi_{i}$ is the Rademacher random variable (taking values ±$\pm 1$ with equal probability $1/2$), $\mathcal{G}_{n}(\mathcal{F};\delta,\xi_{1:n})$ is the local Rademacher complexity and is denoted by $\mathcal{R}_{n}(\mathcal{F};\delta)$.
This definition is a direct generalization of the local Rademacher and Gaussian complexities [Bartlett et al., 2005; Koltchinskii, 2006; Wainwright, 2019]. When $\xi_{1:n}$ is the noise sequence $\eta_{1:n}$, the local complexity $\mathcal{G}_{n}(\mathcal{F};\delta,\eta_{1:n})$ measures how well the function class $\mathcal{F}$ correlates with the noises on the samples $X_{1:n}$. Note that, $\mathcal{G}_{n}(\mathcal{F};\delta,\eta_{1:n})$ is a data-dependent quantity. If the sample points $X_{1:n}$ are random, then $\mathcal{G}_{n}(\mathcal{F};\delta,\eta_{1:n})$ is also a random variable.
To use the local complexity, we often require that the function class $\mathcal{F}$ is star-shaped (around the origin), meaning that, for any $f\in\mathcal{F}$ and $a\in[0,1]$, the function $af\in\mathcal{F}$. If the star-shaped condition fails to hold, one can consider the star hull
$\,{\rm star}\,(\mathcal{F}):=\{af:f\in\mathcal{F},a\in[0,1]\}.$
One of the key properties of the local complexities of a star-shaped class $\mathcal{F}$ is that the function $\delta\mapsto\mathcal{G}_{n}(\mathcal{F};\delta,\xi_{1:n})/\delta$ is non-increasing on $(0,\infty)$. This can be easily proven by observing that the rescaled function $\widetilde{f}=\frac{\delta_{1}}{\delta_{2}}f\in\mathcal{F}$ if $f\in\mathcal{F}$ and $0<\delta_{1}\leq\delta_{2}$ [Wainwright, 2019, Lemma 13.6]. Consequently, for any constant $c>0$, the inequality $\mathcal{G}_{n}(\mathcal{F};\delta,\xi_{1:n})\leq c\delta^{2}$ always has positive solutions $\delta=\delta_{n}(\mathcal{F})$. Our proof of Theorem 3.1 uses these solutions to control the complexity of neural networks $\mathcal{C}_{n}(\mathcal{F}_{n})\lesssim\delta_{n}^{2}(\mathcal{F}_{n})$ (see Lemmas 5.2, 5.3 and 5.5). Note that the function classes $\mathcal{F}_{\sigma}(M)$ and $\mathcal{NN}(N,M)$ are star-shaped. The following theorem gives an estimate for the local complexities of shallow neural networks, which holds uniformly for all sample points $X_{1:n}$ and network width.
Theorem 3.3.
Let $d\geq 4$ and $\xi_{1:n}$ be a sequence of independent sub-Gaussian random variables with parameter $\varsigma>0$ in the sense that
$\mathbb{E}[\exp(\lambda\xi_{i})]\leq\exp(\varsigma^{2}\lambda^{2}/2),\quad\forall\lambda\in\mathbb{R}.$
Then, for any $0<\delta\leq M$,
$\mathcal{G}_{n}(\mathcal{F}_{\sigma}(M);\delta,\xi_{1:n})\lesssim\frac{\varsigma\delta^{\frac{3}{d+3}}M^{\frac{d}{d+3}}}{\sqrt{n}}\sqrt{\log(nM/\delta)},$
where the implied constant is independent of $\xi_{1:n}$ and the sample points $X_{1:n}$ in $\mathbb{B}^{d}$. The bound also holds for the function class $\,{\rm star}\,(\mathcal{T}_{B}\mathcal{F}_{\sigma}(M))$ for any constant $B>0$.
Note that, when $\delta=M$, we obtain the bound $\mathcal{O}(M\sqrt{\log n}/\sqrt{n})$, which is the same as the bound for (global) Rademacher complexity derived in [Golowich et al., 2020, Theorem 3.2], if we ignore the logarithmic factor. The key point of Theorem 3.3 is that the bound is independent of the network width so that we can apply it to over-parameterized neural networks. Since the network width $N_{n}$ dose not affect the complexity $\mathcal{C}_{n}(\mathcal{F}_{n})$, one can simply increase $N_{n}$ in Theorem 2.2 to decrease the approximation error $\mathcal{E}_{app}(\mathcal{H},\mathcal{F}_{n})$ (when $\mathcal{H}=\mathcal{H}^{\alpha}$, the error is dominated by the norm constraint $M_{n}$). Theorem 3.1 gives a lower bound on $N_{n}$ to attain the desired accuracy.
3.2 Regularized least squares
The least squares estimate (3.1) with $\mathcal{F}_{n}=\mathcal{NN}(N_{n},M_{n})$ is a constrained optimization problem, where we make constraint on the norm $\kappa(f)\leq M_{n}$. One can also estimate the regression function $h$ by solving the corresponding regularized optimization problem
$\operatorname*{argmin}_{f\in\mathcal{NN}(N_{n})}\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})-Y_{i})^{2}+\lambda_{n}\kappa(f),$ (3.2)
where $\lambda_{n}>0$ is a tunable parameter. Usually, the regularization parameter $\lambda_{n}$ is chosen to satisfy $\lim_{n\to\infty}\lambda_{n}=0$. We show in the next theorem that, if $\lambda_{n}$ is chosen properly, the regularized least squares has the same rate of convergence as the least squares estimator. The result is also true if the norm $\kappa(f)$ in (3.2) is replaced by $\gamma(f)$.
Theorem 3.4.
Let $d\geq 4$ and $\widehat{f}_{n,\lambda_{n}}$ be a solution of the optimization problem (3.2).
  1. (1)
    If $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$, we choose
    $N_{n}\gtrsim n^{\frac{d}{d+2\alpha}},\quad\lambda_{n}\asymp n^{-\frac{d+3+2\alpha}{2d+4\alpha}}\log n,$
    then the bound
    $\|\mathcal{T}_{B}\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{2\alpha}{d+2\alpha}}\log n$
    holds with probability at least $1-c_{1}\exp(-c_{2}n^{\frac{d}{d+2\alpha}}\log n)$ for some constants $c_{1},c_{2}>0$.
  2. (2)
    If $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, we choose
    $N_{n}\gtrsim n^{\frac{d}{2d+3}},\quad\lambda_{n}\asymp n^{-\frac{d+3}{2d+3}}\log n,$
    then the bound
    $\|\mathcal{T}_{B}\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{d+3}{2d+3}}\log n$
    holds with probability at least $1-c_{1}\exp(-c_{2}n^{\frac{d}{2d+3}}\log n)$ for some constants $c_{1},c_{2}>0$.
Comparing with Theorem 3.1, the main difficulty in proving Theorem 3.4 is that we do not have explicit bound on the norm $\kappa(\widehat{f}_{n,\lambda_{n}})$. In order to apply Theorems 2.2 and 3.3, one need to estimate the norm of the learned function from the optimization problem (3.2). A trivial bound is to choose $f=0$ and get
$\kappa(\widehat{f}_{n,\lambda_{n}})\leq\frac{1}{n\lambda_{n}}\sum_{i=1}^{n}Y_{i}^{2}\leq\frac{B^{2}}{\lambda_{n}}.$
Unfortunately, this bound is always too loose to obtain the optimal rate. To see this, let $f^{*}$ be a minimizer of the regularized approximation error
$\mathcal{E}_{\lambda_{n}}:=\inf_{f\in\mathcal{NN}(N_{n})}\|f-h\|_{L^{2}(\mu)}^{2}+\lambda_{n}\kappa(f),$
which can be regarded as the expected version of the regularized least squares (3.2). Then, we know that $\kappa(f^{*})\leq\mathcal{E}_{\lambda_{n}}/\lambda_{n}$. Yet, we expect $\widehat{f}_{n,\lambda_{n}}$ to be a good approximation of $f^{*}$. So, one would expect that $\kappa(\widehat{f}_{n,\lambda_{n}})\lesssim\mathcal{E}_{\lambda_{n}}/\lambda_{n}$, which is always better than the bound $B^{2}/\lambda_{n}$, because $\mathcal{E}_{\lambda_{n}}\to 0$ if $\lambda_{n}$ is chosen properly. This phenomenon is also discussed in detail by Wu et al. [2005] and Steinwart and Christmann [2008, Chapter 7] in the analysis of support vector machines. They also discussed how to use iteration technique to improve the trivial bound. Nonetheless, we overcome the difficulty by using a new oracle inequality for the regularized least squares estimator (see Lemma 5.5). Our proof shows that one can apply this oracle inequality to obtain a bound close to $\mathcal{E}_{\lambda_{n}}/\lambda_{n}$
In practice, the neural network $f_{\theta}\in\mathcal{NN}(N_{n})$ is parameterized by $\theta$ in the form (2.3). By Theorem 2.1, we can estimate $\kappa(f)$ by $\kappa(\theta)$, which reduces the optimization problem (3.2) to
$\operatorname*{argmin}_{\theta\in\mathbb{R}^{(d+2)N_{n}}}\frac{1}{n}\sum_{i=1}^{n}(f_{\theta}(X_{i})-Y_{i})^{2}+\lambda_{n}\kappa(\theta),$ (3.3)
One can further simplify the optimization problem by observing that the regularizer $\kappa(\theta)\leq\|\theta\|_{2}^{2}/2$. It may be easier to solve the following optimization problem
$\operatorname*{argmin}_{\theta\in\mathbb{R}^{(d+2)N_{n}}}\frac{1}{n}\sum_{i=1}^{n}(f_{\theta}(X_{i})-Y_{i})^{2}+\frac{\lambda_{n}}{2}\|\theta\|_{2}^{2}.$ (3.4)
The next proposition shows that the two optimization problems (3.3) and (3.4) essentially have the same solutions. To the best of our knowledge, this equivalence is first pointed out by [Neyshabur et al., 2015b, Theorem 1].
Proposition 3.5.
If $\theta^{*}$ is a solution of the optimization problem (3.3), then there exits a solution $\widetilde{\theta}^{*}$ of (3.4) such that $f_{\theta^{*}}=f_{\widetilde{\theta}^{*}}$. Any solution of (3.4) is also a solution of (3.3).
Proof.
Let us denote $N=N_{n}$ for simplicity. Due to the positive homogeneity of the ReLU function, we can rescale the parameter $\theta=(a_{1},w_{1}^{\intercal},\dots,a_{N},w_{N}^{\intercal})^{\intercal}$ to $(c_{1}a_{1},c_{1}^{-1}w_{1}^{\intercal},\dots,c_{N}a_{N},c_{N}^{-1}w_{N}^{\intercal})^{\intercal}$ with $c_{i}>0$, without altering the function $f_{\theta}$. We use $r(\theta)=(\widetilde{a}_{1},\widetilde{w}_{1}^{\intercal},\dots,\widetilde{a}_{N},\widetilde{w}_{N}^{\intercal})^{\intercal}$ to denote the rescaled parameter of $\theta$ that satisfies $|\widetilde{a}_{i}|=\|\widetilde{w}_{i}\|_{2}$ for all $1\leq i\leq N$ (if $|a_{i}|\|w_{i}\|_{2}=0$, we let $|\widetilde{a}_{i}|=\|\widetilde{w}_{i}\|_{2}=0$). Note that we always have $\kappa(\theta)=\kappa(r(\theta))=\|r(\theta)\|_{2}^{2}/2\leq\|\theta\|_{2}^{2}/2$.
Let us denote the empirical square loss of $f_{\theta}$ by $\mathcal{L}_{n}(f_{\theta})=\frac{1}{n}\sum_{i=1}^{n}(f_{\theta}(X_{i})-Y_{i})^{2}$ for convenience. If $\theta^{*}$ is a solution of (3.3), then $r(\theta^{*})$ is also a solution of (3.3). Thus, for any $\theta$, we have
$\mathcal{L}_{n}(f_{r(\theta^{*})})+\frac{\lambda_{n}}{2}\|r(\theta^{*})\|_{2}^{2}=\mathcal{L}_{n}(f_{r(\theta^{*})})+\lambda_{n}\kappa(r(\theta^{*}))\leq\mathcal{L}_{n}(f_{\theta})+\lambda_{n}\kappa(\theta)\leq\mathcal{L}_{n}(f_{\theta})+\frac{\lambda_{n}}{2}\|\theta\|_{2}^{2},$
which shows that $r(\theta^{*})$ is a solution of (3.4). Conversely, if $\theta^{*}$ is a solution of (3.4), then $r(\theta^{*})=\theta^{*}$ because $\|r(\theta)\|_{2}^{2}\leq\|\theta\|_{2}^{2}$. Thus, for any $\theta$,
$\mathcal{L}_{n}(f_{\theta^{*}})+\lambda_{n}\kappa(\theta^{*})=\mathcal{L}_{n}(f_{\theta^{*}})+\frac{\lambda_{n}}{2}\|\theta^{*}\|_{2}^{2}\leq\mathcal{L}_{n}(f_{r(\theta)})+\frac{\lambda_{n}}{2}\|r(\theta)\|_{2}^{2}=\mathcal{L}_{n}(f_{\theta})+\kappa(\theta),$
which shows that $\theta^{*}$ also minimizes (3.3). ∎
Using the same proof idea as Theorem 3.4, one can show that the solutions of the regularized optimization problems (3.3) and (3.4) can also achieve the minimax optimal rates for learning functions in $\mathcal{H}^{\alpha}$ or $\mathcal{F}_{\sigma}(1)$.
Corollary 3.6.
Let $d\geq 4$ and $\widehat{f}_{n,\lambda_{n}}=f_{\widehat{\theta}_{n,\lambda_{n}}}$, where $\widehat{\theta}_{n,\lambda_{n}}$ is a solution of the optimization problem (3.3) or (3.4). Then, the conclusion of Theorem 3.4 also holds.
4 Discussions and related works
[…]
There is another line of works [Bartolucci et al., 2023; Ongie et al., 2020; Parhi and Nowak, 2021, 2022, 2023; Savarese et al., 2019] trying to characterize infinitely wide neural networks from a functional analysis point of view. These papers derived representer theorems showing that (finitely wide) neural networks are sparse solutions to data fitting problems with total variation regularization in the Radon domain:
$\operatorname*{argmin}_{f}\frac{1}{n}\sum_{i=1}^{n}\ell(Y_{i},f(X_{i}))+\lambda\,{\rm RTV}\,(f),$ (4.1)
where the loss function $\ell(\cdot,\cdot)$ is convex and coercive in its second argument and $\,{\rm RTV}\,$ is a seminorm defined through the Radon transform. Note that the null space of the seminorm $\,{\rm RTV}\,$ is the space of affine functions [Parhi and Nowak, 2021, Lemma 19]. For $f\in\mathcal{NN}(N)$ in the reduced form (2.4), $\,{\rm RTV}\,(f)$ is exactly the variation $\sum_{k=1}^{K}|c_{k}|$ by [Parhi and Nowak, 2021, Lemma 25], and hence $\kappa(f)\approx\,{\rm RTV}\,(f)+\|\mbox{liner part of }f\|$ by inequality (2.5). Thus, for square loss $\ell$, we can view the optimization problem (4.1) as the problem (3.2) with width $N_{n}=\infty$.
Infinitely wide neural networks are also used in the study of training dynamics of neural networks. The neural tangent kernel theory demonstrates that the evolution of a neural network during gradient descent training can be described by a kernel in the infinite width limit [Allen-Zhu et al., 2019; Du et al., 2019; Jacot et al., 2018; Oymak and Soltanolkotabi, 2020]. The mean field analysis approximates the evolution of the network weights by a gradient flow, defined through a partial differential equation, in the Wasserstein space of probability distributions [Mei et al., 2018, 2019; Sirignano and Spiliopoulos, 2020]. These theories establish the convergences of gradient descent training of over-parameterized neural networks in certain scalings of network weights [Chizat et al., 2019]. It would be interesting to see whether these techniques can be applied to study the optimization problems (3.1) and (3.3).
Nonparametric regression using neural networks. As discussed in the introduction, the rates of convergence of neural network regression estimates have been analyzed by many authors. Minimax optimal rates have been established for sparse deep neural networks [Schmidt-Hieber, 2020; Suzuki, 2019], for fully connected deep neural networks [Kohler and Langer, 2021] and for shallow neural networks [Yang and Zhou, 2023]. It has also been shown that deep neural networks are able to circumvent the curse of dimensionality under certain conditions, for example, the intrinsic dimension is low [Chen et al., 2022; Jiao et al., 2021; Nakada and Imaizumi, 2020] and the regression function has certain hierarchical structure [Bauer and Kohler, 2019; Schmidt-Hieber, 2020; Kohler and Langer, 2021]. In the over-parameterized regime, [Yang and Zhou, 2023] proved nearly optimal rates by using (deep or shallow) neural networks with weight constraints. The paper [Parhi and Nowak, 2023] established the minimax optimal rate for the variation space $\mathcal{F}_{\sigma}$ using infinitely wide shallow neural networks. Our results can be viewed as generalizations of this result to Hölder classes $\mathcal{H}^{\alpha}$ and finitely wide networks.
Interpolation and regularization. The recent progress in deep learning revealed that over-parameterized models can interpolate noisy data and yet generalize well [Belkin et al., 2019a; Zhang et al., 2017]. Motivated by this phenomenon, which is called benign overfitting, there has recently been a line of works on understanding theoretical mechanisms for the good generalization performance of interpolating models (see [Bartlett et al., 2021; Belkin, 2021] for reviews). In particular, benign overfitting has been theoretically established for many models, such as Nadaraya-Watson smoothing estimator [Belkin et al., 2019b], linear regression [Bartlett et al., 2020; Hastie et al., 2022], kernel regression [Liang and Rakhlin, 2020] and random features models [Mei et al., 2022]. However, this certainly does not mean that any interpolator can generalize well, simply because there are too many estimators that interpolate observed data. Hence, explicit or implicit regularization is necessary to obtain estimators with good generalization.
In fact, most works on benign overfitting studied the minimal norm interpolations. In our setting, this is corresponding to the solutions of
$\operatorname*{argmin}_{f\in\mathcal{NN}(N_{n})}\kappa(f),\quad s.t.\quad f(X_{i})=Y_{i},\ i=1,\dots,n,$ (4.2)
which can be viewed as the limiting estimator of (3.2) with $\lambda_{n}=0$. The setting of over-parameterized neural networks allows us to make an explicit comparison between interpolating and regularized solutions. We think it is a promising research direction to study the rate of convergence for the minimal norm interpolation (4.2) and compare it with our results. Note that, for linear regression, Hastie et al. [2022] showed that optimally regularized ridge regression dominates the minimum $\ell^{2}$-norm interpolation in test performance. It is reasonable to expect that such phenomenon also occurs for neural networks (since the regularized least squares (3.2) achieves the minimax rates). On the other hand, some kernel interpolations have been proved to be inconsistent in low dimensions [Buchholz, 2022; Rakhlin and Zhai, 2019]. It is an interesting problem to determine whether the minimal norm solutions of (4.2) are consistent and minimax optimal.
5 Proofs
This section gives the proofs of Theorems 2.1, 3.1, 3.3, 3.4 and Corollary 3.6.
5.1 Proof of Theorem 2.1
Recall that $S_{-}$ is defined by (2.2), $S_{+}=-S_{-}$ and $S_{0}=\mathbb{S}^{d}\setminus(S_{-}\cup S_{+})$. Notice that $S_{-}$ is the set of vectors $v\in\mathbb{S}^{d}$ such that $(x^{\intercal},1)v\leq 0$ for all $x\in\mathbb{B}^{d}$. Thus, $v\in\mathbb{S}^{d}$ is in $S_{+}$ if and only if $(x^{\intercal},1)v\geq 0$ holds for all $x\in\mathbb{B}^{d}$. And, for any $v\in S_{0}$, $\sigma((x^{\intercal},1)v)$ is a nonlinear function on $\mathbb{B}^{d}$. Let us first characterize the measures $\tau\in\mathcal{M}_{disc}(\mathbb{S}^{d})$ such that $f_{\tau}=0$.
Lemma 5.1.
Let $\tau=\sum_{i=1}^{N}a_{i}\delta_{u_{i}}\in\mathcal{M}_{disc}(\mathbb{S}^{d})$, where $u_{1},\dots,u_{N}\in\mathbb{S}^{d}$ are distinct. Then, $f_{\tau}(x)=\sum_{i=1}^{N}a_{i}\sigma((x^{\intercal},1)u_{i})=0$ for all $x\in\mathbb{B}^{d}$ if and only if the following two conditions hold
  1. (1)
    For any $u_{i}\in S_{0}$, there exists $j\neq i$ such that $u_{j}=-u_{i}$ and $a_{j}=-a_{i}$.
  2. (2)
    $\frac{1}{2}\sum_{u_{i}\in S_{0}}a_{i}u_{i}+\sum_{u_{i}\in S_{+}}a_{i}u_{i}=0$.
Proof.
Without loss of generality, we can assume that $u_{i}\in S_{0}\cup S_{+}$ and $a_{i}\neq 0$ for $i=1,\dots,N$. Notice that $a_{i}(x^{\intercal},1)u_{i}=a_{i}\sigma((x^{\intercal},1)u_{i})-a_{i}\sigma(-(x^{\intercal},1)u_{i})$. If $\tau$ satisfies conditions (1) and (2), then
$f_{\tau}(x)$ $=\sum_{u_{i}\in S_{0}}a_{i}\sigma((x^{\intercal},1)u_{i})+\sum_{u_{i}\in S_{+}}a_{i}(x^{\intercal},1)u_{i}$
$=\frac{1}{2}\sum_{u_{i}\in S_{0}}a_{i}(x^{\intercal},1)u_{i}+\sum_{u_{i}\in S_{+}}a_{i}(x^{\intercal},1)u_{i}=0.$
Conversely, assume $f_{\tau}=0$ on $\mathbb{B}^{d}$. Notice that $\sigma((x^{\intercal},1)u_{i})$ is a piecewise linear function. We denote its “break points” by $A_{i}=\{x:\|x\|_{2}<1,(x^{\intercal},1)u_{i}=0\}$. One can check that $A_{i}$ is non-empty if and only if $u_{i}\in S_{0}$. For any $u_{i}\in S_{0}$, let $j$ be the index such that $u_{j}=-u_{i}$ (it may not exist), then we can find a point $x_{i}\in A_{i}=A_{j}$ such that $x_{i}\notin A_{k}$ for any $k\neq i,j$. Consider the decomposition
$0=f_{\tau}(x)=a_{i}\sigma((x^{\intercal},1)u_{i})+a_{j}\sigma((x^{\intercal},1)u_{j})+\sum_{k\neq i,j}a_{k}\sigma((x^{\intercal},1)u_{k}).$
If $u_{j}$ does not exist, then the first term is nonlinear in a neighborhood of $x_{i}$ and the remain terms are linear in this neighborhood, which is a contradiction. Hence, $u_{j}$ exists and the summation of the first two terms is linear, which implies $a_{j}=-a_{i}$. As a consequence,
$0=f_{\tau}(x)=\frac{1}{2}\sum_{u_{i}\in S_{0}}a_{i}(x^{\intercal},1)u_{i}+\sum_{u_{i}\in S_{+}}a_{i}(x^{\intercal},1)u_{i}=(x^{\intercal},1)\left(\frac{1}{2}\sum_{u_{i}\in S_{0}}a_{i}u_{i}+\sum_{u_{i}\in S_{+}}a_{i}u_{i}\right)$
holds for all $x\in\mathbb{B}^{d}$, Thus, $\frac{1}{2}\sum_{u_{i}\in S_{0}}a_{i}u_{i}+\sum_{u_{i}\in S_{+}}a_{i}u_{i}=0$. ∎
For any $f\in\mathcal{NN}(N)$, we can always represent it as
$f(x)=\sum_{\widetilde{v}_{i}\in S_{0}}\widetilde{c}_{i}\sigma((x^{\intercal},1)\widetilde{v}_{i})+\sum_{\widetilde{v}_{i}\in S_{+}}\widetilde{c}_{i}(x^{\intercal},1)\widetilde{v}_{i}.$ (5.1)
Without loss of generality, we assume that the $\widetilde{v}_{i}$ are distinct. If there exist $\widetilde{v}_{i}$ and $\widetilde{v}_{j}$ satisfying $\widetilde{v}_{j}=-\widetilde{v}_{i}\in S_{0}$, then
$\widetilde{c}_{i}\sigma((x^{\intercal},1)\widetilde{v}_{i})+\widetilde{c}_{j}\sigma((x^{\intercal},1)\widetilde{v}_{j})=(\widetilde{c}_{i}+\widetilde{c}_{j})\sigma((x^{\intercal},1)\widetilde{v}_{i})-\widetilde{c}_{j}(x^{\intercal},1)\widetilde{v}_{i}.$
Using this equality, we can reduce (5.1) to the following form
$f(x)=\sum_{k=1}^{K}c_{k}\sigma((x^{\intercal},1)v_{k})+(x^{\intercal},1)w,$ (5.2)
where $0\leq K\leq N$, $c_{1},\dots,c_{K}\neq 0$, $w\in\mathbb{R}^{d+1}$ and $v_{1},\dots,v_{K}\in S_{0}$ satisfy $v_{i}\neq\pm v_{k}$ for $1\leq i\neq k\leq K$. Note that, if $K=N$, then $w=0$. We are going to construct a measure $\mu\in\mathcal{M}_{disc}(\mathbb{S}^{d})$ supported on at most $N$ points in $S_{0}\cup S_{+}$ such that $f_{\mu}=f$ and estimate $\kappa(f)$ by $\|\mu\|$. By Lemma 5.1, any measure $\nu$ satisfies $f_{\nu}=f_{\mu}=f$ if and only if $\nu=\mu+\tau$ for some $\tau$ satisfying the two conditions in Lemma 5.1. Hence, $\kappa(f)=\inf_{\tau}\|\mu+\tau\|$. Without loss of generality, we also assume that $\tau$ is supported on $S_{0}\cup S_{+}$. By Lemma 5.1, we can decompose $\tau$ as
$\tau=\sum_{k=1}^{K}a_{k}\left(\delta_{v_{k}}-\delta_{-v_{k}}\right)+\sum_{j=1}^{J_{1}}b_{j}\delta_{u_{j}}+\sum_{j=J_{1}+1}^{J}b_{j}(\delta_{u_{j}}-\delta_{-u_{j}})=:\tau_{1}+\tau_{2}+\tau_{3},$
where $u_{j}\in S_{+}$ for $1\leq j\leq J_{1}$ and $u_{j}\in S_{0}\setminus\{\pm v_{1},\dots,\pm v_{K}\}$ for $J_{1}+1\leq j\leq J$ satisfy $u_{j}\neq\pm u_{i}$ for any $i\neq j$ and
$\sum_{k=1}^{K}a_{k}v_{k}+\sum_{j=1}^{J}b_{j}u_{j}=0.$ (5.3)
We consider four cases:
Case 1: $w=0$. We let $\mu=\sum_{k=1}^{K}c_{k}\delta_{v_{k}}$, then
$\|\mu+\tau\|$ $=\|\mu+\tau_{1}\|+\|\tau_{2}+\tau_{3}\|\geq\|\mu+\tau_{1}\|$
$=\sum_{k=1}^{K}\left(|c_{k}+a_{k}|+|a_{k}|\right)\geq\sum_{k=1}^{K}|c_{k}|=\|\mu\|,$
which implies $\kappa(f)=\|\mu\|$.
Case 2: $w/\|w\|_{2}\in S_{+}\cup S_{-}$. We let $\mu=\sum_{k=1}^{K}c_{k}\delta_{v_{k}}+c_{K+1}\delta_{v_{K+1}}=:\mu_{1}+\mu_{2}$, where, if $w/\|w\|_{2}\in S_{+}$, we take $v_{K+1}=w/\|w\|_{2}\in S_{+}$ and $c_{K+1}=\|w\|_{2}$; otherwise, we take $v_{K+1}=-w/\|w\|_{2}\in S_{+}$ and $c_{K+1}=-\|w\|_{2}$. Since $w\neq 0$, $K\leq N-1$ and $\mu$ is supported on at most $K+1\leq N$ points. Let $m\in\{1,\dots,J_{1}\}$ be the index such that $u_{m}=v_{K+1}$. (If such $m$ does not exists, we set $m=0$ and $b_{m}=0$ in the following.) Similar to Case 1, we have
$\|\mu+\tau\|$ $=\|\mu_{1}+\tau_{1}\|+\|\mu_{2}+b_{m}\delta_{u_{m}}\|+\|\tau_{2}-b_{m}\delta_{u_{m}}+\tau_{3}\|$
$\geq\sum_{k=1}^{K}\left(|c_{k}+a_{k}|+|a_{k}|\right)+|c_{K+1}+b_{m}|+\sum_{j\neq m}|b_{j}|.$
By equality (5.3),
$w=c_{K+1}v_{K+1}=\sum_{k=1}^{K}a_{k}v_{k}+(c_{K+1}+b_{m})u_{m}+\sum_{j\neq m}b_{j}u_{j},$
which implies
$\|\mu+\tau\|\geq\sum_{k=1}^{K}|a_{k}|+|c_{K+1}+b_{m}|+\sum_{j\neq m}|b_{j}|\geq\|w\|_{2}.$
Combining with the bound $\|\mu+\tau\|\geq\sum_{k=1}^{K}|c_{k}|$, we conclude that
$\kappa(f)=\inf_{\tau}\|\mu+\tau\|\geq\|w\|_{2}\lor\sum_{k=1}^{K}|c_{k}|.$ (5.4)
Therefore,
$\kappa(f)\geq\frac{1}{2}\sum_{k=1}^{K}|c_{k}|+\frac{1}{2}\|w\|_{2}=\frac{1}{2}\|\mu\|.$
Case 3: $w=cv_{i}$ for some $1\leq i\leq K$ and $c\neq 0$. We let $\mu=\sum_{k=1}^{K}c_{k}\delta_{v_{k}}+c(\delta_{v_{i}}-\delta_{-v_{i}})$, which is supported on at most $K+1\leq N$ points, then $\|\mu\|=\sum_{k\neq i}|c_{k}|+|c_{i}+c|+|c|$ and
$\|\mu+\tau\|$ $=\|\mu+\tau_{1}\|+\|\tau_{2}+\tau_{3}\|$
$\geq\sum_{k\neq i}\left(|c_{k}+a_{k}|+|a_{k}|\right)+|c_{i}+c+a_{i}|+|c+a_{i}|+\sum_{j=1}^{J}|b_{j}|.$
Similar to Case 2, by equality (5.3),
$w=cv_{i}=\sum_{k\neq i}a_{k}v_{k}+(c+a_{i})v_{i}+\sum_{j=1}^{J}b_{j}u_{j},$
which implies
$\|\mu+\tau\|\geq\sum_{k\neq i}|a_{k}|+|c+a_{i}|+\sum_{j=1}^{J}|b_{j}|\geq\|w\|_{2}.$
Combining with $\|\mu+\tau\|\geq\sum_{k=1}^{K}|c_{k}|$, we can obtain the bound (5.4). Since $|c|=\|w\|_{2}$,
$\|\mu\|=\sum_{k\neq i}|c_{k}|+|c_{i}+c|+|c|\leq\sum_{k=1}^{K}|c_{k}|+2\|w\|_{2}\leq 3\kappa(f).$
Case 4: $v_{K+1}:=w/\|w\|_{2}\in S_{0}\setminus\{\pm v_{1},\dots,\pm v_{K}\}$. By the construction of (5.2), it holds that $K\leq N-2$. We let $\mu=\sum_{k=1}^{K}c_{k}\delta_{v_{k}}+\|w\|_{2}(\delta_{v_{K+1}}-\delta_{-v_{K+1}})=:\mu_{1}+\mu_{2}$, which is supported on $K+2\leq N$ points. Let $m\in\{J_{1}+1,\dots,J\}$ be the index such that $u_{m}=\pm v_{K+1}$. (If such $m$ does not exists, we set $m=0$ and $b_{m}=0$ in the following.) Then,
$\|\mu+\tau\|$ $=\|\mu_{1}+\tau_{1}\|+\|\mu_{2}+b_{m}(\delta_{u_{m}}-\delta_{-u_{m}})\|+\|\tau_{2}+\tau_{3}-b_{m}(\delta_{u_{m}}-\delta_{-u_{m}})\|$
$\geq\sum_{k=1}^{K}\left(|c_{k}+a_{k}|+|a_{k}|\right)+|\|w\|_{2}\pm b_{m}|+\sum_{j\neq m}|b_{j}|.$
Similar to Case 2, by using equality (5.3), we can obtain the bound (5.4). Therefore,
$\|\mu\|=\sum_{k=1}^{K}|c_{k}|+2\|w\|_{2}\leq 3\kappa(f).$
In summary, we have shown that, for the parameterization (5.2), there exists a measure $\mu$ supported on at most $N$ points such that $f_{\mu}=f$ and
$\|w\|_{2}\lor\sum_{k=1}^{K}|c_{k}|\leq\kappa(f)\leq\|\mu\|\leq\sum_{k=1}^{K}|c_{k}|+2\|w\|_{2}.$
It is easy to construct $\theta\in\mathbb{R}^{(d+2)N}$ such that $f_{\mu}=f_{\theta}$ and $\kappa(\theta)=\|\mu\|$. Hence, $\kappa(f)\leq\inf\{\kappa(\theta):f=f_{\theta},\theta\in\mathbb{R}^{(d+2)N}\}\leq 3\kappa(f)$.
5.2 Proof of Theorem 3.1
We use similar idea as Nakada and Imaizumi [2020]; Schmidt-Hieber [2020] (see also Wainwright [2019, Chapters 13 and 14]). We first give a decomposition of the error. For any $f\in\mathcal{F}_{n}$ (we will take $f$ to minimize the approximation error $\|f-h\|_{L^{\infty}(\mathbb{B}^{d})})$, observe that
$\|\mathcal{T}_{B}\widehat{f}_{n}-h\|_{L^{2}(\mu)}^{2}$ $\leq 2\|\mathcal{T}_{B}\widehat{f}_{n}-\mathcal{T}_{B}f\|_{L^{2}(\mu)}^{2}+2\|\mathcal{T}_{B}f-h\|_{L^{2}(\mu)}^{2}$
$\leq 2\|\mathcal{T}_{2B}(\widehat{f}_{n}-f)\|_{L^{2}(\mu)}^{2}+2\|f-h\|_{L^{2}(\mu)}^{2},$ (5.5)
where we use $|\mathcal{T}_{B}\widehat{f}_{n}(x)-\mathcal{T}_{B}f(x)|\leq|\mathcal{T}_{2B}(\widehat{f}_{n}-f)(x)|$ and $\|h\|_{L^{\infty}(\mathbb{B}^{d})}\leq B$ in the second inequality. Note that, for the empirical error,
$\|\mathcal{T}_{2B}(\widehat{f}_{n}-f)\|_{L^{2}(\mu_{n})}^{2}\leq\|\widehat{f}_{n}-f\|_{L^{2}(\mu_{n})}^{2}\leq 2\|\widehat{f}_{n}-h\|_{L^{2}(\mu_{n})}^{2}+2\|f-h\|_{L^{2}(\mu_{n})}^{2}.$ (5.6)
By the definition of $\widehat{f}_{n}$,
$\frac{1}{n}\sum_{i=1}^{n}(\widehat{f}_{n}(X_{i})-Y_{i})^{2}\leq\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})-Y_{i})^{2}.$
Using $Y_{i}=h(X_{i})+\eta_{i}$, we have the base inequality
$\|\widehat{f}_{n}-h\|_{L^{2}(\mu_{n})}^{2}\leq\|f-h\|_{L^{2}(\mu_{n})}^{2}+\frac{2}{n}\sum_{i=1}^{n}\eta_{i}\left(\widehat{f}_{n}(X_{i})-f(X_{i})\right)$ (5.7)
Our proof can be divided into following three steps.
  1. Step 1.
    Estimating $\|\widehat{f}_{n}-h\|_{L^{2}(\mu_{n})}^{2}$ by using the base inequality (5.7).
  2. Step 2.
    Bounding $\|\mathcal{T}_{2B}(\widehat{f}_{n}-f)\|_{L^{2}(\mu)}^{2}$ by its empirical counterpart $\|\mathcal{T}_{2B}(\widehat{f}_{n}-f)\|_{L^{2}(\mu_{n})}^{2}$.
  3. Step 3.
    Combining inequalities (5.5), (5.6), Step 1, Step 2 and the approximation result (Theorem 2.2).
We now give the detail of the proof. We will use $c_{1},c_{2}$ to denote constants, which may be different in different bounds.
Step 1. We derive an oracle inequality for the least squares estimator as Wainwright [2019, Theorem 13.13], which assumes that the noise is standard Gaussian. Our proof is essentially the same as the proof of [Wainwright, 2019, Theorem 13.13]. Since we will use similar argument in the proof of Lemma 5.5, we give the proof here for completeness. It is convenient to use the notation :=$\partial\mathcal{F}:=\{f_{1}-f_{2}:f_{1},f_{2}\in\mathcal{F}\}$.
Lemma 5.2.
For any fixed sample points $X_{1:n}$, let $\widehat{f}_{n}$ be the estimator (3.1) and $\delta_{n}$ be any positive solution to the inequality
$\mathcal{G}_{n}(\,{\rm star}\,(\partial\mathcal{F}_{n});\delta_{n},\eta_{1:n})\leq\delta_{n}^{2}.$
Then, there are constants $c_{1},c_{2}>0$ such that the bound
$\|\widehat{f}_{n}-h\|_{L^{2}(\mu_{n})}^{2}\leq 3\|f-h\|_{L^{2}(\mu_{n})}^{2}+32\delta_{n}^{2},\quad\forall f\in\mathcal{F}_{n},$
holds with probability at least $1-c_{1}\exp(-c_{2}n\delta_{n}^{2}/B^{2})$.
Proof.
If $f\in\mathcal{F}_{n}$ satisfies $\|\widehat{f}_{n}-f\|_{L^{2}(\mu_{n})}\leq\delta_{n}$, then
$\|\widehat{f}_{n}-h\|_{L^{2}(\mu_{n})}^{2}$ $\leq 2\|f-h\|_{L^{2}(\mu_{n})}^{2}+2\|\widehat{f}_{n}-f\|_{L^{2}(\mu_{n})}^{2}$
$\leq 2\|f-h\|_{L^{2}(\mu_{n})}^{2}+2\delta_{n}^{2}.$
Thus, we can assume $\|\widehat{f}_{n}-f\|_{L^{2}(\mu_{n})}>\delta_{n}$.
Observing that $\widehat{f}_{n}-f\in\partial\mathcal{F}_{n}$, we consider the event
$\mathcal{A}:=\left\{\exists g\in\partial\mathcal{F}_{n}:\|g\|_{L^{2}(\mu_{n})}>\delta_{n},\left|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}g(X_{i})\right|>2\delta_{n}\|g\|_{L^{2}(\mu_{n})}\right\}.$
If $\mathcal{A}$ is true, then there exists $\widetilde{g}:=\frac{\delta_{n}}{\|g\|_{L^{2}(\mu_{n})}}g\in\,{\rm star}\,(\partial\mathcal{F}_{n})$ such that $\|\widetilde{g}\|_{L^{2}(\mu_{n})}=\delta_{n}$ and
$\left|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}\widetilde{g}(X_{i})\right|=\frac{\delta_{n}}{\|g\|_{L^{2}(\mu_{n})}}\left|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}g(X_{i})\right|>2\delta_{n}^{2}.$
This implies that $\mathbb{P}(\mathcal{A})\leq\mathbb{P}(Z_{n}>2\delta_{n}^{2})$, where $Z_{n}$ is the random variable defined by
$Z_{n}:=\sup_{\begin{subarray}{c}g\in\,{\rm star}\,(\partial\mathcal{F}_{n})\\ \|g\|_{L^{2}(\mu_{n})}\leq\delta_{n}\end{subarray}}\left|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}g(X_{i})\right|.$
Observe that we can view $Z_{n}(\eta_{1:n})$ as a convex Lipschitz function of the random vector $\eta_{1:n}$ supported on $[-2B,2B]^{n}$ with Lipschitz constant $\delta_{n}/\sqrt{n}$. By the concentration for convex Lipschitz functions on bounded random variables [Boucheron et al., 2013, Theorem 6.10], we conclude that, for any $t>0$,
$\mathbb{P}(Z_{n}>\mathbb{E}[Z_{n}]+t)\leq\exp\left(-\frac{nt^{2}}{32B^{2}\delta_{n}^{2}}\right).$ (5.8)
Taking $t=\delta_{n}^{2}$ and noticing that $\mathbb{E}[Z_{n}]=\mathcal{G}_{n}(\,{\rm star}\,(\partial\mathcal{F}_{n});\delta_{n},\eta_{1:n})\leq\delta_{n}^{2}$ by assumption, we get
$\mathbb{P}(\mathcal{A})\leq\mathbb{P}(Z_{n}>2\delta_{n}^{2})\leq\mathbb{P}(Z_{n}>\mathbb{E}[Z_{n}]+\delta_{n}^{2})\leq\exp\left(-\frac{n\delta_{n}^{2}}{32B^{2}}\right).$
By the definition of the event $\mathcal{A}$ and the assumption $\|\widehat{f}_{n}-f\|_{L^{2}(\mu_{n})}>\delta_{n}$, the inequality
$\left|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}\left(\widehat{f}_{n}(X_{i})-f(X_{i})\right)\right|\leq 2\delta_{n}\|\widehat{f}_{n}-f\|_{L^{2}(\mu_{n})}$
holds with probability at least $1-\exp(-n\delta_{n}^{2}/(32B^{2}))$. Combining with the base inequality (5.7), we have
$\|\widehat{f}_{n}-h\|_{L^{2}(\mu_{n})}^{2}$ $\leq\|f-h\|_{L^{2}(\mu_{n})}^{2}+4\delta_{n}\|\widehat{f}_{n}-f\|_{L^{2}(\mu_{n})}$
$\leq\|f-h\|_{L^{2}(\mu_{n})}^{2}+16\delta_{n}^{2}+\frac{1}{4}\|\widehat{f}_{n}-f\|_{L^{2}(\mu_{n})}^{2}$
$\leq\|f-h\|_{L^{2}(\mu_{n})}^{2}+16\delta_{n}^{2}+\frac{1}{2}\|\widehat{f}_{n}-h\|_{L^{2}(\mu_{n})}^{2}+\frac{1}{2}\|f-h\|_{L^{2}(\mu_{n})}^{2}$
$=\frac{3}{2}\|f-h\|_{L^{2}(\mu_{n})}^{2}+16\delta_{n}^{2}+\frac{1}{2}\|\widehat{f}_{n}-h\|_{L^{2}(\mu_{n})}^{2}.$
Rearranging yields the desired bound. ∎
We apply Lemma 5.2 to $\mathcal{F}_{n}=\mathcal{NN}(N_{n},M_{n})$. Since $\,{\rm star}\,(\partial\mathcal{F}_{n})\subseteq\mathcal{F}_{\sigma}(2M_{n})$, by Theorem 3.3, we have
$\mathcal{G}_{n}(\,{\rm star}\,(\partial\mathcal{F}_{n});\delta_{n},\eta_{1:n})\lesssim\frac{B\delta_{n}^{\frac{3}{d+3}}M_{n}^{\frac{d}{d+3}}}{\sqrt{n}}\sqrt{\log(nM_{n}/\delta_{n})},$
for all sample points $X_{1:n}$ in $\mathbb{B}^{d}$. Hence, we can choose
$\delta_{n}\asymp n^{-\frac{d+3}{4d+6}}M_{n}^{\frac{d}{2d+3}}\sqrt{\log(nM_{n})}.$
Then, with probability at least $1-c_{1}\exp(-c_{2}n^{\frac{d}{2d+3}}M_{n}^{\frac{2d}{2d+3}}\log(nM_{n}))$, it holds
$\|\widehat{f}_{n}-h\|_{L^{2}(\mu_{n})}^{2}\lesssim\|f-h\|_{L^{2}(\mu_{n})}^{2}+n^{-\frac{d+3}{2d+3}}M_{n}^{\frac{2d}{2d+3}}\log(nM_{n}),\quad\forall f\in\mathcal{F}_{n}.$ (5.9)
Step 2. To bound $\|\mathcal{T}_{2B}(\widehat{f}_{n}-f)\|_{L^{2}(\mu)}^{2}$, we use the following uniform law with localization from [Wainwright, 2019, Theorem 14.1 and Proposition 14.25]. In the lemma, it is required that the function class $\mathcal{F}$ is star-shaped and $B$-uniformly bounded for some constant $B>0$, meaning that $\|f\|_{L^{\infty}(\mathbb{B}^{d})}\leq B$ for all $f\in\mathcal{F}$.
Lemma 5.3.
Given a star-shaped and $B$-uniformly bounded function class $\mathcal{F}$, let $\epsilon_{n}$ be any positive solution to the inequality
$\mathcal{R}_{n}(\mathcal{F};\epsilon_{n})\leq\frac{\epsilon_{n}^{2}}{B}.$
Then, there are constants $c_{1},c_{2}>0$ such that the bound
$\|f\|_{L^{2}(\mu)}^{2}\leq 2\|f\|_{L^{2}(\mu_{n})}^{2}+\epsilon_{n}^{2},\quad\forall f\in\mathcal{F},$
holds with probability at least $1-c_{1}\exp(-c_{2}n\epsilon_{n}^{2}/B^{2})$.
Observe that
$\mathcal{T}_{2B}(\widehat{f}_{n}-f)\in\mathcal{T}_{2B}\partial\mathcal{F}_{n}\subseteq\mathcal{T}_{2B}\mathcal{F}_{\sigma}(2M_{n})\subseteq\,{\rm star}\,(\mathcal{T}_{2B}\mathcal{F}_{\sigma}(2M_{n})).$
We can apply Lemma 5.3 to the star-shaped and $2B$-uniformly bounded function class $\,{\rm star}\,(\mathcal{T}_{2B}\mathcal{F}_{\sigma}(2M_{n}))$. By Theorem 3.3,
$\mathcal{R}_{n}(\,{\rm star}\,(\mathcal{T}_{2B}\mathcal{F}_{\sigma}(2M_{n}));\epsilon_{n})\lesssim\frac{\epsilon_{n}^{\frac{3}{d+3}}M_{n}^{\frac{d}{d+3}}}{\sqrt{n}}\sqrt{\log(nM_{n}/\epsilon_{n})},$
and hence we can choose
$\epsilon_{n}\asymp n^{-\frac{d+3}{4d+6}}M_{n}^{\frac{d}{2d+3}}\sqrt{\log(nM_{n})}.$
Then, with probability at least $1-c_{1}\exp(-c_{2}n^{\frac{d}{2d+3}}M_{n}^{\frac{2d}{2d+3}}\log(nM_{n}))$, it holds
$\|\mathcal{T}_{2B}(\widehat{f}_{n}-f)\|_{L^{2}(\mu)}^{2}\lesssim\|\mathcal{T}_{2B}(\widehat{f}_{n}-f)\|_{L^{2}(\mu_{n})}^{2}+n^{-\frac{d+3}{2d+3}}M_{n}^{\frac{2d}{2d+3}}\log(nM_{n}),\quad\forall f\in\mathcal{F}_{n}.$ (5.10)
Step 3. Combining inequalities (5.5), (5.6), (5.9) and (5.10) and , we get
$\|\mathcal{T}_{B}\widehat{f}_{n}-h\|_{L^{2}(\mu)}^{2}\lesssim\|f-h\|_{L^{2}(\mu)}^{2}+\|f-h\|_{L^{2}(\mu_{n})}^{2}+n^{-\frac{d+3}{2d+3}}M_{n}^{\frac{2d}{2d+3}}\log(nM_{n}),$
which holds with probability at least $1-c_{1}\exp(-c_{2}n^{\frac{d}{2d+3}}M_{n}^{\frac{2d}{2d+3}}\log(nM_{n}))$ simultaneously for all $f\in\mathcal{F}_{n}$. Taking infimum over $f\in\mathcal{F}_{n}$ shows
$\|\mathcal{T}_{B}\widehat{f}_{n}-h\|_{L^{2}(\mu)}^{2}\lesssim\inf_{f\in\mathcal{F}_{n}}\|f-h\|_{L^{\infty}(\mathbb{B}^{d})}^{2}+n^{-\frac{d+3}{2d+3}}M_{n}^{\frac{2d}{2d+3}}\log(nM_{n}).$
(1) If $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$, then by Theorem 2.2,
$\sup_{h\in\mathcal{H}}\inf_{f\in\mathcal{F}_{n}}\|f-h\|_{L^{\infty}(\mathbb{B}^{d})}^{2}\lesssim N_{n}^{-\frac{2\alpha}{d}}\lor M_{n}^{-\frac{4\alpha}{d+3-2\alpha}}.$
If $M_{n}\asymp n^{\frac{d+3-2\alpha}{2d+4\alpha}}$ and $N_{n}\gtrsim n^{\frac{d}{d+2\alpha}}$, then
$\|\mathcal{T}_{B}\widehat{f}_{n}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{2\alpha}{d+2\alpha}}\log n,$
holds with probability at least $1-c_{1}\exp(-c_{2}n^{\frac{d}{d+2\alpha}}\log n)$.
(2) If $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, then by Theorem 2.2, there exist a constant $M\geq 1$ such that for any $M_{n}=M$,
$\sup_{h\in\mathcal{H}}\inf_{f\in\mathcal{F}_{n}}\|f-h\|_{L^{\infty}(\mathbb{B}^{d})}^{2}\lesssim N_{n}^{-\frac{d+3}{d}}.$
If $N_{n}\gtrsim n^{\frac{d}{2d+3}}$, then
$\|\mathcal{T}_{B}\widehat{f}_{n}-h\|_{L^{2}(\mu)}^{2}\lesssim n^{-\frac{d+3}{2d+3}}\log n,$
holds with probability at least $1-c_{1}\exp(-c_{2}n^{\frac{d}{2d+3}}\log n)$.
Remark 5.4.
The boundedness of the noises $\eta_{1:n}$ is used to derive the dimension-free high probability bound (5.8) (Note that the $n$ in the exponential is from the Lipschitz constant). The proof relies on the concentration inequality for convex Lipschitz functions. This argument can also be applied to the sub-Gaussian noises characterized by Gozlan et al. [2017, Corollary 5.11] (which of course includes Gaussian). Hence, similar results as Theorem 3.1 can be obtained for such kind of noises.
5.3 Proof of Theorem 3.3
Without loss of generality, we can assume that $\varsigma=1$ by rescaling. Note that the sub-Gaussian assumption $\mathbb{E}[\exp(\lambda\xi_{i})]\leq\exp(\lambda^{2}/2)$ implies $\mathbb{E}[\xi_{i}]=0$ and $\mathbb{E}[\xi_{i}^{2}]\leq 1$. It is well known that the Gaussian complexity can be bounded by Dudley’s entropy integral [Wainwright, 2019, Section 5.3.3]. We will provide a bound for the entropy by using Theorem 2.2.
To be concrete, for any $\epsilon>0$, let us denote the $\epsilon$-covering number of the set $\mathcal{F}$ in $L^{2}(\mu_{n})$ norm by $\mathcal{N}(\epsilon,\mathcal{F},\|\cdot\|_{L^{2}(\mu_{n})})$. In other words, there exists a subset $\{f_{1},\dots,f_{m}\}\subseteq\mathcal{F}$ of minimal size $m=\mathcal{N}(\epsilon,\mathcal{F},\|\cdot\|_{L^{2}(\mu_{n})})$ such that for any $f\in\mathcal{F}$, there exits $f_{j}$ satisfying $\|f-f_{j}\|_{L^{2}(\mu_{n})}\leq\epsilon$. Then,
$\left|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}f(X_{i})\right|$ $\leq\left|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}f_{j}(X_{i})\right|+\left|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}(f(X_{i})-f_{j}(X_{i}))\right|$
$\leq\max_{j=1,\dots,m}\left|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}f_{j}(X_{i})\right|+\sqrt{\frac{\sum_{i=1}^{n}\xi_{i}^{2}}{n}}\epsilon,$
where we use the Cauchy-Schwarz inequality in the second step. Taking the supremum over $f\in\mathcal{F}$ and then expectation over the noises, we obtain
$\mathbb{E}_{\xi_{1:n}}\left[\sup_{f\in\mathcal{F}}\left|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}f(X_{i})\right|\right]\leq\mathbb{E}_{\xi_{1:n}}\left[\max_{j=1,\dots,m}\left|\frac{1}{n}\sum_{i=1}^{n}\xi_{i}f_{j}(X_{i})\right|\right]+\epsilon,$ (5.11)
where we have used $\mathbb{E}[(\sum_{i=1}^{n}\xi_{i}^{2})^{1/2}]\leq(\sum_{i=1}^{n}\mathbb{E}[\xi_{i}^{2}])^{1/2}\leq\sqrt{n}$. Define a family of zero-mean random variables indexed by $f\in\mathcal{F}$ as $Z(f):=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\xi_{i}f(X_{i})$. Then, $Z(f)$ is a sub-Gaussian process with respect to the metric $\rho_{Z}(f,f^{\prime}):=\|f-f^{\prime}\|_{L^{2}(\mu_{n})}$, meaning that
$\mathbb{E}\left[\exp(\lambda(Z(f)-Z(f^{\prime})))\right]\leq\exp(\lambda^{2}\rho_{Z}(f,f^{\prime})^{2}/2),\quad\forall f,f^{\prime}\in\mathcal{F},\lambda\in\mathbb{R},$
which can be proven by basic properties of sub-Gaussian variables [Wainwright, 2019, Exercise 2.13]. Observe that the first term on the right-hand side of (5.11) is an expected supremum of the sub-Gaussian process $Z(f)$. We can apply a chaining argument as Wainwright [2019, Proof of Theorem 5.22] to show that
[…]
where $D:=\sup_{f,f^{\prime}\in\mathcal{F}}\rho_{Z}(f,f^{\prime})$ denotes the diameter. Therefore,
[…]
In order to apply this bound to the variation space $\mathcal{F}_{\sigma}(M)$, we need to estimate its covering number. Since $\|f-f^{\prime}\|_{L^{2}(\mu_{n})}\leq\|f-f^{\prime}\|_{L^{\infty}(\mathbb{B}^{d})}$, it is enough to provide a bound for the entropy $\log\mathcal{N}(\epsilon,\mathcal{F}_{\sigma}(M),\|\cdot\|_{L^{\infty}(\mathbb{B}^{d})})$. Note that this bound should be independent of the network width. But we will need to first bound the entropy of $\mathcal{NN}(N,M)$ by $N$ and $M$. Recall that any $f\in\mathcal{NN}(N,M)$ can be represented in the form
$f(x)=\sum_{i=1}^{N}a_{i}\sigma((x^{\intercal},1)v_{i}),\quad v_{i}\in\mathbb{S}^{d},\ \sum_{i=1}^{N}|a_{i}|\leq M.$
It is easy to see that there exists $\epsilon/(2\sqrt{2}M)$-cover $V_{\epsilon}$ of $\mathbb{S}^{d}$ in the metric $\|\cdot\|_{2}$ with cardinality $|V_{\epsilon}|\lesssim(1+M/\epsilon)^{d}$. Similarly, there exists $\epsilon/(2\sqrt{2})$-cover $A_{\epsilon}$ of the set $A=\{a=(a_{1},\dots,a_{N}):\|a\|_{1}\leq M\}$ in the metric $\|\cdot\|_{1}$ such that $|A_{\epsilon}|\lesssim(1+M/\epsilon)^{N}$. Hence, for any $v_{i}\in\mathbb{S}^{d}$ and $a=(a_{1},\dots,a_{N})\in A$, we can choose $\widetilde{v}_{i}\in V_{\epsilon}$ and $\widetilde{a}=(\widetilde{a}_{1},\dots,\widetilde{a}_{N})\in A_{\epsilon}$ such that $\|v_{i}-\widetilde{v}_{i}\|_{2}\leq\epsilon/(2\sqrt{2}M)$ and $\|a-\widetilde{a}\|_{1}\leq\epsilon/(2\sqrt{2})$. Therefore, letting $\widetilde{f}(x)=\sum_{i=1}^{N}\widetilde{a}_{i}\sigma((x^{\intercal},1)\widetilde{v}_{i})$, we have for $x\in\mathbb{B}^{d}$,
$|f(x)-\widetilde{f}(x)|$ $\leq\|a-\widetilde{a}\|_{1}\sup_{1\leq i\leq N}|\sigma((x^{\intercal},1)v_{i})|+\|\widetilde{a}\|_{1}\sup_{1\leq i\leq N}|\sigma((x^{\intercal},1)v_{i})-\sigma((x^{\intercal},1)\widetilde{v}_{i})|$
$\leq\epsilon/2+\sqrt{2}M\sup_{1\leq i\leq N}\|v_{i}-\widetilde{v}_{i}\|_{2}\leq\epsilon,$
where we use the Lipschitz continuity of ReLU in the second inequality. Thus,
$\log\mathcal{N}(\epsilon,\mathcal{NN}(N,M),\|\cdot\|_{L^{\infty}(\mathbb{B}^{d})})\lesssim N\log(1+M/\epsilon).$ (5.13)
Now, we estimate the entropy of $\mathcal{F}_{\sigma}(M)$. By Theorem 2.2, there exists $M_{0}>0$ such that
$\sup_{f\in\mathcal{F}_{\sigma}(M)}\inf_{f_{N}\in\mathcal{NN}(N,M_{0}M)}\|f-f_{N}\|_{L^{\infty}(\mathbb{B}^{d})}\lesssim MN^{-\frac{1}{2}-\frac{3}{2d}}.$
Hence, for any $0<\epsilon\leq M$, we can choose $N\asymp(\epsilon/M)^{-2d/(d+3)}$ such that for any $f\in\mathcal{F}_{\sigma}(M)$, there exists $f_{N}\in\mathcal{NN}(N,M_{0}M)$ satisfying $\|f-f_{N}\|_{L^{\infty}(\mathbb{B}^{d})}<\epsilon/2$. By the triangle inequality, any $\epsilon/2$-cover of $\mathcal{NN}(N,M_{0}M)$ is an $\epsilon$-cover of $\mathcal{F}_{\sigma}(M)$. Then, the estimate (5.13) shows
$\log\mathcal{N}(\epsilon,\mathcal{F}_{\sigma}(M),\|\cdot\|_{L^{\infty}(\mathbb{B}^{d})})$ $\leq\log\mathcal{N}(\epsilon/2,\mathcal{NN}(N,M_{0}M),\|\cdot\|_{L^{\infty}(\mathbb{B}^{d})})$
$\lesssim N\log(1+2M_{0}M/\epsilon)$
$\lesssim(\epsilon/M)^{-2d/(d+3)}\log(1+M/\epsilon).$
Applying (5.12) to =$\mathcal{F}=\{f\in\mathcal{F}_{\sigma}(M):\|f\|_{L^{2}(\mu_{n})}\leq\delta\}$, we have $D=2\delta$ and
[…]
If we choose $\epsilon\asymp\delta^{\frac{3}{d+3}}M^{\frac{d}{d+3}}n^{-1/2}$, then for $n\geq 2$,
[…]
$\lesssim\frac{\delta^{\frac{3}{d+3}}M^{\frac{d}{d+3}}}{\sqrt{n}}\sqrt{\log(nM/\delta)}.$
Finally, for the star hull $\,{\rm star}\,(\mathcal{T}_{B}\mathcal{F}_{\sigma}(M))$, we can similarly derive the bound by estimating the covering number. It is easy to see that the covering number of $\mathcal{T}_{B}\mathcal{F}_{\sigma}(M)$ is not larger than that of $\mathcal{F}_{\sigma}(M)$. For any $\epsilon>0$, let $\{f_{1},\dots,f_{m}\}\subseteq\mathcal{T}_{B}\mathcal{F}_{\sigma}(M)$ be an $\epsilon/2$-cover of $\mathcal{T}_{B}\mathcal{F}_{\sigma}(M)$ and $\{a_{1},\dots,a_{k}\}\subseteq[0,1]$ be an $\epsilon/(2B)$-cover of $[0,1]$. Then, for any $af\in\,{\rm star}\,(\mathcal{T}_{B}\mathcal{F}_{\sigma}(M))$, we can choose $a_{i}f_{j}$ such that
$\|af-a_{i}f_{j}\|_{L^{\infty}(\mathbb{B}^{d})}\leq|a-a_{i}|\|f\|_{L^{\infty}(\mathbb{B}^{d})}+|a_{i}|\|f-f_{j}\|_{L^{\infty}(\mathbb{B}^{d})}\leq\epsilon.$
This shows that $\{a_{i}f_{j}:1\leq i\leq k,1\leq j\leq m\}$ is $\epsilon$-cover of $\,{\rm star}\,(\mathcal{T}_{B}\mathcal{F}_{\sigma}(M))$, which implies
$\log\mathcal{N}(\epsilon,\,{\rm star}\,(\mathcal{T}_{B}\mathcal{F}_{\sigma}(M)),\|\cdot\|_{L^{\infty}(\mathbb{B}^{d})})$
$\leq$ $\log\mathcal{N}(\epsilon,\mathcal{F}_{\sigma}(M),\|\cdot\|_{L^{\infty}(\mathbb{B}^{d})})+\log(1+B/\epsilon)$
$\lesssim$ $(\epsilon/M)^{-2d/(d+3)}\log(1+M/\epsilon).$
Hence, we can obtain the same bound for the local complexity.
5.4 Proof of Theorem 3.4
As in the proof of Theorem 3.1, for any $f\in\mathcal{NN}(N_{n})$, we have
$\|\mathcal{T}_{B}\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu)}^{2}$ $\leq 2\|\mathcal{T}_{2B}(\widehat{f}_{n,\lambda_{n}}-f)\|_{L^{2}(\mu)}^{2}+2\|f-h\|_{L^{2}(\mu)}^{2},$ (5.14)
$\|\mathcal{T}_{2B}(\widehat{f}_{n,\lambda_{n}}-f)\|_{L^{2}(\mu_{n})}^{2}$ $\leq 2\|\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu_{n})}^{2}+2\|f-h\|_{L^{2}(\mu_{n})}^{2}.$ (5.15)
By the definition of $\widehat{f}_{n,\lambda_{n}}$,
$\frac{1}{n}\sum_{i=1}^{n}(\widehat{f}_{n,\lambda_{n}}(X_{i})-Y_{i})^{2}+\lambda_{n}\kappa(\widehat{f}_{n,\lambda_{n}})\leq\frac{1}{n}\sum_{i=1}^{n}(f(X_{i})-Y_{i})^{2}+\lambda_{n}\kappa(f).$
Using $Y_{i}=h(X_{i})+\eta_{i}$, we get the base inequality
$\|\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(\widehat{f}_{n,\lambda_{n}})\leq\|f-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(f)+\frac{2}{n}\sum_{i=1}^{n}\eta_{i}\left(\widehat{f}_{n,\lambda_{n}}(X_{i})-f(X_{i})\right).$ (5.16)
Similar to Theorem 3.1, our proof strategy contains three parts:
  1. 1.
    Estimating $\|\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu_{n})}^{2}$ and $\kappa(\widehat{f}_{n,\lambda_{n}})$ by using the base inequality (5.16).
  2. 2.
    Estimating $\|\mathcal{T}_{2B}(\widehat{f}_{n,\lambda_{n}}-f)\|_{L^{2}(\mu)}^{2}$ by its empirical counterpart $\|\mathcal{T}_{2B}(\widehat{f}_{n,\lambda_{n}}-f)\|_{L^{2}(\mu_{n})}^{2}$.
  3. 3.
    Combining the estimates with inequalities (5.14), (5.15) and the approximation result.
The last two parts can be done in a similar way as Theorem 3.1. For the first part, we use the following new oracle inequality, which is a modification of Lemma 5.2 to regularized estimators. Recall that :=$\partial\mathcal{F}:=\{f_{1}-f_{2}:f_{1},f_{2}\in\mathcal{F}\}$.
Lemma 5.5.
For any fixed sample points $X_{1:n}$ and constant $c_{0}\geq 1$, let $\widehat{f}_{n,\lambda_{n}}\in\mathcal{NN}(N_{n})$ be any estimator (i.e. a function depending on the random noises $\eta_{1:n}$, where $\mathbb{E}[\eta_{i}]=0$ and $|\eta_{i}|\leq 2B$) that satisfies the inequality
$\|\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(\widehat{f}_{n,\lambda_{n}})$ (5.17)
$\leq c_{0}\left(\|f-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(f)\right)+\frac{2}{n}\left|\sum_{i=1}^{n}\eta_{i}\left(\widehat{f}_{n,\lambda_{n}}(X_{i})-f(X_{i})\right)\right|,\quad\forall f\in\mathcal{NN}(N_{n}).$
For a user-defined parameter $R>0$, let $\delta_{n}=\delta_{n}(R)$ be any positive solution to the inequality
$\mathcal{G}_{n}(\,{\rm star}\,(\partial\mathcal{F}_{n,R});\delta_{n},\eta_{1:n})\leq\delta_{n}^{2},$
where $\mathcal{F}_{n,R}:=\{f\in\mathcal{NN}(N_{n}):\kappa(f)\leq R\}$. There are constants $c_{1},c_{2}>0$ such that, if $\lambda_{n}\geq 8\delta_{n}^{2}/R$, then the bound
$\|\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(\widehat{f}_{n,\lambda_{n}})\leq(1+2c_{0})\left(\|f-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(f)\right)+64\delta_{n}^{2},\quad\forall f\in\mathcal{F}_{n,R},$
holds with probability at least $1-c_{1}\exp(-c_{2}n\delta_{n}^{2}/B^{2})$.
To apply Lemma 5.5, let us first compute $\delta_{n}=\delta_{n}(R)$, where $R>0$ will be chosen later. Since $\,{\rm star}\,(\partial\mathcal{F}_{n,R})\subseteq\mathcal{F}_{\sigma}(2R)$, by Theorem 3.3,
$\mathcal{G}_{n}(\,{\rm star}\,(\partial\mathcal{F}_{n,R});\delta_{n},\eta_{1:n})\lesssim\frac{B\delta_{n}^{\frac{3}{d+3}}R^{\frac{d}{d+3}}}{\sqrt{n}}\sqrt{\log(nR/\delta_{n})},$
holds for all sample points $X_{1:n}$ in $\mathbb{B}^{d}$. Hence, we can choose
$\delta_{n}^{2}\asymp n^{-\frac{d+3}{2d+3}}R^{\frac{2d}{2d+3}}\log(nR).$
We denote the approximation error of $\mathcal{F}_{n,R}$ by
$\mathcal{E}(R):=\inf_{f\in\mathcal{F}_{n,R}}\|f-h\|_{L^{\infty}(\mathbb{B}^{d})}^{2}.$
Note that we use supremum norm instead of $L^{2}(\mu_{n})$ norm in the definition. Lemma 5.5 with $c_{0}=1$ shows that, if $\lambda_{n}=8\delta_{n}^{2}/R$, then with probability at least $1-c_{1}\exp(-c_{2}n\delta_{n}^{2})$, it holds
$\|\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(\widehat{f}_{n,\lambda_{n}})\lesssim\mathcal{E}(R)+\delta_{n}^{2},$
In particular, we have
$\kappa(\widehat{f}_{n,\lambda_{n}})\lesssim\frac{\mathcal{E}(R)+\delta_{n}^{2}}{\lambda_{n}}\lesssim\frac{\mathcal{E}(R)}{\lambda_{n}}+R,$
where we have used $\delta_{n}^{2}\lesssim\lambda_{n}R$ by the choice of $\lambda_{n}$.
(1) If $\mathcal{H}=\mathcal{H}^{\alpha}$ with $\alpha<(d+3)/2$, then by Theorem 2.2,
$\mathcal{E}(R)\leq\sup_{h\in\mathcal{H}}\inf_{f\in\mathcal{F}_{n,R}}\|f-h\|_{L^{\infty}(\mathbb{B}^{d})}^{2}\lesssim N_{n}^{-\frac{2\alpha}{d}}\lor R^{-\frac{4\alpha}{d+3-2\alpha}}.$
We choose
$R\asymp n^{\frac{d+3-2\alpha}{2d+4\alpha}},\quad N_{n}\gtrsim R^{\frac{2d}{d+3-2\alpha}}\gtrsim n^{\frac{d}{d+2\alpha}},$
then
$\mathcal{E}(R)\lesssim n^{-\frac{2\alpha}{d+2\alpha}},\quad\delta_{n}^{2}\asymp n^{-\frac{2\alpha}{d+2\alpha}}\log n,\quad\lambda_{n}=6\delta_{n}^{2}/R\asymp n^{-\frac{d+3+2\alpha}{2d+4\alpha}}\log n.$ (5.18)
(2) If $\mathcal{H}=\mathcal{F}_{\sigma}(1)$, then by Theorem 2.2, there exist a constant $M\geq 1$ such that for any $R\geq M$,
$\mathcal{E}(R)\leq\sup_{h\in\mathcal{H}}\inf_{f\in\mathcal{F}_{n,R}}\|f-h\|_{L^{\infty}(\mathbb{B}^{d})}^{2}\lesssim N_{n}^{-\frac{d+3}{d}}.$
We choose
$R=M,\quad N_{n}\gtrsim n^{\frac{d}{2d+3}},$
then
$\mathcal{E}(R)\lesssim n^{-\frac{d+3}{2d+3}},\quad\delta_{n}^{2}\asymp n^{-\frac{d+3}{2d+3}}\log n,\quad\lambda_{n}=6\delta_{n}^{2}/R\asymp n^{-\frac{d+3}{2d+3}}\log n.$ (5.19)
Notice that, in both cases, we always have $\mathcal{E}(R)\lesssim\lambda_{n}R$, which implies $M_{n}:=\kappa(\widehat{f}_{n,\lambda_{n}})\lesssim R$. Similar to step 2 in the proof of Theorem 3.1, we can apply Lemma 5.3 to the function class $\,{\rm star}\,(\mathcal{T}_{2B}\mathcal{F}_{\sigma}(M_{n}+R))$ to show that
$\|\mathcal{T}_{2B}(\widehat{f}_{n,\lambda_{n}}-f)\|_{L^{2}(\mu)}^{2}\lesssim\|\mathcal{T}_{2B}(\widehat{f}_{n,\lambda_{n}}-f)\|_{L^{2}(\mu_{n})}^{2}+\epsilon_{n}^{2},\quad\forall f\in\mathcal{F}_{n,R},$
holds with probability at least $1-c_{1}\exp(-c_{2}n\epsilon_{n}^{2})$, where
$\epsilon_{n}^{2}\asymp n^{-\frac{d+3}{2d+3}}R^{\frac{2d}{2d+3}}\log(nR)\asymp\delta_{n}^{2}.$
Combining with inequalities (5.14) and (5.15), where we choose $f\in\mathcal{F}_{n,R}$ to satisfy the bound $\|f-h\|_{L^{\infty}(\mathbb{B}^{d})}^{2}\lesssim\mathcal{E}(R)$, we obtain
$\|\mathcal{T}_{B}\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu)}^{2}\lesssim\mathcal{E}(R)+\delta_{n}^{2},$
holds with probability at least $1-c_{1}\exp(-c_{2}n\delta_{n}^{2})$. The estimates (5.18) and (5.19) then give the desired result.
To complete the proof of Theorem 3.4, it remains to prove Lemma 5.5.
Proof of Lemma 5.5.
We modify the proofs of Lemma 5.2 and [Wainwright, 2019, Theorem 13.17]. For any $f\in\mathcal{F}_{n,R}$, let us denote $g_{f}:=\widehat{f}_{n,\lambda_{n}}-f$ for convenience. We divide the proof into four cases.
Case 1: $\|g_{f}\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(g_{f})\leq\delta_{n}^{2}$. We have
$\|\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(\widehat{f}_{n,\lambda_{n}})$
$\leq$ $2\|f-h\|_{L^{2}(\mu_{n})}^{2}+2\|g_{f}\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(g_{f})+\lambda_{n}\kappa(f)$
$\leq$ $2\|f-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(f)+2\delta_{n}^{2}.$
Case 2: $\|g_{f}\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(g_{f})>\delta_{n}^{2}$ and $\kappa(\widehat{f}_{n,\lambda_{n}})\leq R$. In this case, $g_{f}=\widehat{f}_{n,\lambda_{n}}-f\in\partial\mathcal{F}_{n,R}$. Consider the event
$\mathcal{A}_{1}:=\left\{\exists g\in\partial\mathcal{F}_{n,R}:\|g\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(g)>\delta_{n}^{2},\left|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}g(X_{i})\right|>2\delta_{n}\sqrt{\|g\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(g)}\right\}.$
If $\mathcal{A}_{1}$ is true, then there exists
$\widetilde{g}:=\frac{\delta_{n}}{\sqrt{\|g\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(g)}}g\in\,{\rm star}\,(\partial\mathcal{F}_{n,R}),$
such that $\|\widetilde{g}\|_{L^{2}(\mu_{n})}\leq\delta_{n}$ and
$\left|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}\widetilde{g}(X_{i})\right|=\frac{\delta_{n}}{\sqrt{\|g\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(g)}}\left|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}g(X_{i})\right|>2\delta_{n}^{2}.$
This implies that $\mathbb{P}(\mathcal{A}_{1})\leq\mathbb{P}(Z_{n}(\delta_{n})>2\delta_{n}^{2})$, where $Z_{n}(\delta_{n})$ is the random variable defined by
$Z_{n}(\delta_{n}):=\sup_{\begin{subarray}{c}g\in\,{\rm star}\,(\partial\mathcal{F}_{n,R})\\ \|g\|_{L^{2}(\mu_{n})}\leq\delta_{n}\end{subarray}}\left|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}g(X_{i})\right|.$
As in the proof of Lemma 5.2, by viewing $Z_{n}(\delta_{n})$ as a convex Lipschitz function of $\eta_{1:n}$, we get, for any $u>0$,
$\mathbb{P}(Z_{n}(\delta_{n})>\mathbb{E}[Z_{n}(\delta_{n})]+u)\leq\exp\left(-\frac{nu^{2}}{32B^{2}\delta_{n}^{2}}\right).$
Taking $u=\delta_{n}^{2}$ and noticing that $\mathbb{E}[Z_{n}(\delta_{n})]=\mathcal{G}_{n}(\,{\rm star}\,(\partial\mathcal{F}_{n,R};\delta_{n},\eta_{1:n})\leq\delta_{n}^{2}$, we conclude
$\mathbb{P}(\mathcal{A}_{1})\leq\mathbb{P}(Z_{n}(\delta_{n})>2\delta_{n}^{2})\leq\mathbb{P}(Z_{n}(\delta_{n})>\mathbb{E}[Z_{n}(\delta_{n})]+\delta_{n}^{2})\leq\exp\left(-\frac{n\delta_{n}^{2}}{32B^{2}}\right).$
This implies that the inequality
$\left|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}g_{f}(X_{i})\right|\leq 2\delta_{n}\sqrt{\|g_{f}\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(g_{f})}$
holds with probability at least $1-\exp(-n\delta_{n}^{2}/(32B^{2}))$. Combining with inequality (5.17), we have
$\|\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(\widehat{f}_{n,\lambda_{n}})$
$\leq$ $c_{0}\left(\|f-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(f)\right)+4\delta_{n}\sqrt{\|g_{f}\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(g_{f})}$
$\leq$ $c_{0}\left(\|f-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(f)\right)+16\delta_{n}^{2}+\frac{1}{4}\|\widehat{f}_{n,\lambda_{n}}-f\|_{L^{2}(\mu_{n})}^{2}+\frac{1}{4}\lambda_{n}\kappa(\widehat{f}_{n,\lambda_{n}}-f)$
$\leq$ $c_{0}\left(\|f-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(f)\right)+16\delta_{n}^{2}+\frac{1}{2}\|\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu_{n})}^{2}+\frac{1}{2}\|f-h\|_{L^{2}(\mu_{n})}^{2}$
$\qquad+\frac{1}{4}\lambda_{n}\kappa(\widehat{f}_{n,\lambda_{n}})+\frac{1}{4}\lambda_{n}\kappa(f)$
$\leq$ $\frac{1+2c_{0}}{2}\left(\|f-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(f)\right)+16\delta_{n}^{2}+\frac{1}{2}\left(\|\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(\widehat{f}_{n,\lambda_{n}})\right).$
Rearranging shows that the desired bound holds with probability at least $1-\exp(-n\delta_{n}^{2}/(32B^{2}))$.
Case 3: $\kappa(\widehat{f}_{n,\lambda_{n}})>R$ and $\|g_{f}\|_{L^{2}(\mu_{n})}\leq\delta_{n}\kappa(\widehat{f}_{n,\lambda_{n}})/R$. Observe that
$\widetilde{g}_{f}:=\frac{R}{\kappa(\widehat{f}_{n,\lambda_{n}})}g_{f}=\frac{R}{\kappa(\widehat{f}_{n,\lambda_{n}})}(\widehat{f}_{n,\lambda_{n}}-f)\in\partial\mathcal{F}_{n,R},$
$\|\widetilde{g}_{f}\|_{L^{2}(\mu_{n})}=\frac{R}{\kappa(\widehat{f}_{n,\lambda_{n}})}\|g_{f}\|_{L^{2}(\mu_{n})}\leq\delta_{n}.$
We have shown in Case 2 that $\mathbb{P}(Z_{n}(\delta_{n})>2\delta_{n}^{2})\leq\exp(-n\delta_{n}^{2}/(32B^{2}))$. This implies that the inequality
$\left|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}g_{f}(X_{i})\right|=\frac{\kappa(\widehat{f}_{n,\lambda_{n}})}{R}\left|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}\widetilde{g}_{f}(X_{i})\right|\leq 2\delta_{n}^{2}\frac{\kappa(\widehat{f}_{n,\lambda_{n}})}{R}$
holds with probability at least $1-\exp(-n\delta_{n}^{2}/(32B^{2}))$. By inequality (5.17),
$\|\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(\widehat{f}_{n,\lambda_{n}})$ $\leq c_{0}\left(\|f-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(f)\right)+\frac{4\delta_{n}^{2}}{R}\kappa(\widehat{f}_{n,\lambda_{n}})$
$\leq c_{0}\left(\|f-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(f)\right)+\frac{1}{2}\lambda_{n}\kappa(\widehat{f}_{n,\lambda_{n}}),$
where we use the assumption $\lambda_{n}\geq 8\delta_{n}^{2}/R$ in the last inequality. Therefore,
$\|\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(\widehat{f}_{n,\lambda_{n}})\leq 2c_{0}\left(\|f-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(f)\right),$
holds with probability at least $1-\exp(-n\delta_{n}^{2}/(32B^{2}))$.
Case 4: $\kappa(\widehat{f}_{n,\lambda_{n}})>R$ and $\|g_{f}\|_{L^{2}(\mu_{n})}>\delta_{n}\kappa(\widehat{f}_{n,\lambda_{n}})/R$. In this case,
$\widetilde{g}_{f}=\frac{R}{\kappa(\widehat{f}_{n,\lambda_{n}})}g_{f}\in\partial\mathcal{F}_{n,R},\qquad\|\widetilde{g}_{f}\|_{L^{2}(\mu_{n})}>\delta_{n}.$
We are going to show that the event
$\mathcal{A}_{2}:=\left\{\exists g\in\partial\mathcal{F}_{n,R}:\|g\|_{L^{2}(\mu_{n})}>\delta_{n},\left|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}g(X_{i})\right|>2\delta_{n}\|g\|_{L^{2}(\mu_{n})}+\frac{1}{16}\|g\|_{L^{2}(\mu_{n})}^{2}\right\}$
holds with small probability. We will prove it by a “peeling” argument. To do this, let us denote $t_{m}=2^{m}\delta_{n}$ for $m=0,1,\dots$, and
$\mathcal{A}_{2}(t_{m}):=\left\{\exists g\in\partial\mathcal{F}_{n,R}:t_{m}<\|g\|_{L^{2}(\mu_{n})}\leq 2t_{m},\left|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}g(X_{i})\right|>2\delta_{n}\|g\|_{L^{2}(\mu_{n})}+\frac{1}{16}\|g\|_{L^{2}(\mu_{n})}^{2}\right\},$
then we have the decomposition $\mathcal{A}_{2}=\cup_{m=0}^{\infty}\mathcal{A}_{2}(t_{m})$. To estimate $\mathbb{P}(\mathcal{A}_{2}(t_{m}))$, we consider the random variables
$Z_{n}(t):=\sup_{\begin{subarray}{c}g\in\,{\rm star}\,(\partial\mathcal{F}_{n,R})\\ \|g\|_{L^{2}(\mu_{n})}\leq t\end{subarray}}\left|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}g(X_{i})\right|,\quad t\geq\delta_{n}.$
Notice that, if the event $\mathcal{A}_{2}(t_{m})$ holds, then there exits some function $g\in\partial\mathcal{F}_{n,R}$ with $\|g\|_{L^{2}(\mu_{n})}\in(t_{m},t_{m+1}]$ such that
$\left|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}g(X_{i})\right|$ $>2\delta_{n}\|g\|_{L^{2}(\mu_{n})}+\frac{1}{16}\|g\|_{L^{2}(\mu_{n})}^{2}$
$\geq 2t_{m}\delta_{n}+\frac{1}{16}t_{m}^{2}=t_{m+1}\delta_{n}+\frac{1}{64}t_{m+1}^{2},$
where we use $t_{m+1}=2t_{m}$. This lower bound implies
$\mathbb{P}(\mathcal{A}_{2}(t_{m}))\leq\mathbb{P}\left(Z_{n}(t_{m+1})>t_{m+1}\delta_{n}+\frac{1}{64}t_{m+1}^{2}\right).$
As in Case 2, by viewing $Z_{n}(t)$ as a convex Lipschitz function of $\eta_{1:n}$, one obtain
$\mathbb{P}(Z_{n}(t)>\mathbb{E}[Z_{n}(t)]+u)\leq\exp\left(-\frac{nu^{2}}{32B^{2}t^{2}}\right),\quad\forall u>0.$
Since $\,{\rm star}\,(\partial\mathcal{F}_{n,R})$ is star-shaped, the function $\delta\mapsto\mathcal{G}_{n}(\,{\rm star}\,(\partial\mathcal{F}_{n,R});\delta,\eta_{1:n})/\delta$ is non-increasing. Hence, for any $t\geq\delta_{n}$,
$\mathbb{E}[Z_{n}(t)]=\mathcal{G}_{n}(\,{\rm star}\,(\partial\mathcal{F}_{n,R});t,\eta_{1:n})\leq t\frac{\mathcal{G}_{n}(\,{\rm star}\,(\partial\mathcal{F}_{n,R});\delta_{n},\eta_{1:n})}{\delta_{n}}\leq t\delta_{n}.$
Using this upper bound on the mean and setting $u=t^{2}/64$, we get
$\mathbb{P}\left(Z_{n}(t)>t\delta_{n}+\frac{1}{64}t^{2}\right)\leq\exp\left(-\frac{nt^{2}}{2^{17}B^{2}}\right).$
As a consequence,
$\mathbb{P}(\mathcal{A}_{2})$ $\leq\sum_{m=0}^{\infty}\mathbb{P}(\mathcal{A}_{2}(t_{m}))\leq\sum_{m=1}^{\infty}\mathbb{P}\left(Z_{n}(t_{m})>t_{m}\delta_{n}+\frac{1}{64}t_{m}^{2}\right)$
$\leq\sum_{m=1}^{\infty}\exp\left(-\frac{nt_{m}^{2}}{2^{17}B^{2}}\right)=\sum_{m=1}^{\infty}\exp\left(-\frac{2^{2m}n\delta_{n}^{2}}{2^{17}B^{2}}\right)\leq c_{1}\exp\left(-\frac{c_{2}n\delta_{n}^{2}}{B^{2}}\right),$
for some constant $c_{1},c_{2}>0$.
By the definition of the event $\mathcal{A}_{2}$ and the assumption that $\widetilde{g}_{f}\in\partial\mathcal{F}_{n,R}$ and $\|\widetilde{g}_{f}\|_{L^{2}(\mu_{n})}>\delta_{n}$, we have shown
$\left|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}\widetilde{g}_{f}(X_{i})\right|\leq 2\delta_{n}\|\widetilde{g}_{f}\|_{L^{2}(\mu_{n})}+\frac{1}{16}\|\widetilde{g}_{f}\|_{L^{2}(\mu_{n})}^{2},$
holds with probability at least $1-c_{1}\exp(-c_{2}n\delta_{n}^{2}/B^{2})$. Multiplying both sides by $\kappa(\widehat{f}_{n,\lambda_{n}})/R$, we obtain
$\left|\frac{1}{n}\sum_{i=1}^{n}\eta_{i}g_{f}(X_{i})\right|$ $\leq 2\delta_{n}\|g_{f}\|_{L^{2}(\mu_{n})}+\frac{1}{16}\frac{R}{\kappa(\widehat{f}_{n,\lambda_{n}})}\|g_{f}\|_{L^{2}(\mu_{n})}^{2}$
$\leq 2\delta_{n}\|g_{f}\|_{L^{2}(\mu_{n})}+\frac{1}{16}\|g_{f}\|_{L^{2}(\mu_{n})}^{2},$
where we use $\kappa(\widehat{f}_{n,\lambda_{n}})>R$ in the last inequality. Combining this bound with inequality (5.17), we have
$\|\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(\widehat{f}_{n,\lambda_{n}})$
$\leq$ $c_{0}\left(\|f-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(f)\right)+4\delta_{n}\|g_{f}\|_{L^{2}(\mu_{n})}+\frac{1}{8}\|g_{f}\|_{L^{2}(\mu_{n})}^{2}$
$\leq$ $c_{0}\left(\|f-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(f)\right)+32\delta_{n}^{2}+\frac{1}{4}\|g_{f}\|_{L^{2}(\mu_{n})}^{2}$
$\leq$ $c_{0}\left(\|f-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(f)\right)+32\delta_{n}^{2}+\frac{1}{2}\|\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu_{n})}^{2}+\frac{1}{2}\|f-h\|_{L^{2}(\mu_{n})}^{2}$
$\leq$ $\frac{1+2c_{0}}{2}\left(\|f-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(f)\right)+32\delta_{n}^{2}+\frac{1}{2}\left(\|\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(\widehat{f}_{n,\lambda_{n}})\right),$
which yields the desired bound by rearrangement. ∎
5.5 Proof of Corollary 3.6
By Proposition 3.5, it is enough to consider the case that $\widehat{\theta}_{n,\lambda_{n}}$ is a solution of the optimization problem (3.3). As in the proof of Theorem 3.4, the inequalities (5.14) and (5.15) still hold for $\widehat{f}_{n,\lambda_{n}}$. The only difference is the base inequality
$\|\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(\widehat{\theta}_{n,\lambda_{n}})\leq\|f-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(\theta)+\frac{2}{n}\sum_{i=1}^{n}\eta_{i}\left(\widehat{f}_{n,\lambda_{n}}(X_{i})-f(X_{i})\right),$
where $\theta\in\mathbb{R}^{(d+2)N_{n}}$ satisfy $f_{\theta}=f\in\mathcal{NN}(N_{n})$. By Theorem 2.1, we can choose a $\theta$ such that $\kappa(\theta)\leq 3\kappa(f)$. Hence,
$\|\widehat{f}_{n,\lambda_{n}}-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(\widehat{f}_{n,\lambda_{n}})\leq 3\left(\|f-h\|_{L^{2}(\mu_{n})}^{2}+\lambda_{n}\kappa(f)\right)+\frac{2}{n}\sum_{i=1}^{n}\eta_{i}\left(\widehat{f}_{n,\lambda_{n}}(X_{i})-f(X_{i})\right),$
where we also use $\kappa(\widehat{f}_{n,\lambda_{n}})\leq\kappa(\widehat{\theta}_{n,\lambda_{n}})$. Therefore, we can apply Lemma 5.5 with $c_{0}=3$. The remained proof is the same as Theorem 3.4.
Acknowledgments
The work described in this paper was partially supported by InnoHK initiative, The Government of the HKSAR, Laboratory for AI-Powered Financial Technologies, the Research Grants Council of Hong Kong [Projects No. CityU 11308121, 11306220, 11308020 and No. N_CityU102/20] and National Natural Science Foundation of China [Projects No. 11461161006].